### v1

Abstract
Randomization tests yield exact finite-sample Type 1 error control when the null satisfies the randomization hypothesis. However, achieving these guarantees in practice often requires stronger conditions than the null hypothesis of primary interest. For instance, sign-change tests for mean zero require symmetry and fail to control finite-sample error for non-symmetric mean-zero distributions. We investigate whether such limitations stem from specific test choices or reflect a fundamental inability to construct valid randomization tests for certain hypotheses. We develop a framework providing a simple necessary and sufficient condition for when null hypotheses admit randomization tests. Applying this framework to one-sample tests, we provide characterizations of which nulls satisfy this condition for both finite and continuous supports. In doing so, we prove that certain null hypotheses — including mean zero — do not admit randomization tests. We further show that nulls that admit randomization tests based on linear group actions correspond only to subsets of symmetric or normal distributions. Overall, our findings affirm that practitioners are not inadvertently incurring additional Type 1 error when using existing tests and further motivate focusing on the asymptotic validity of randomization tests.
1 Introduction
Consider the problem of testing a null given a finite sample of data. If this null satisfies a group invariance property referred to as the “randomization hypothesis,” then one can construct randomization tests that obtain exact finite-sample Type 1 error control. Intuitively, this property requires that, under the null, there are transformations of the data that preserve the distribution of the data. In part due to the allure of finite-sample error control, there is a large methodological literature on randomization tests and a similarly large applied literature that makes use of these tests (for a review, see Ritzwoller et al. 2024).
Randomization tests that are used in practice often require stronger conditions than the null hypothesis of primary interest. This is a problem because it re-introduces an inability to control Type 1 error: we may over-reject when the data is drawn from a distribution that satisfies the hypothesis of primary interest but not the stronger conditions. For instance, a commonly-used randomization test for the null of mean zero given i.i.d. observations is constructed by repeatedly recomputing a test statistic under all possible sign changes on the observations. However, this test only properly controls finite-sample error when the data is drawn from a symmetric distribution, and thus may over-reject when data is drawn from a non-symmetric distribution with mean zero.
This paper investigates whether the failure to control Type 1 error for certain null hypotheses (e.g. the null of mean zero) is due to the specific randomization tests used in practice (e.g. the use of sign changes) or due to a more general inability to construct randomization tests that are valid for the hypotheses. To formalize this distinction, we develop a framework to characterize which null hypotheses can and cannot be tested using a randomization test. In doing so, we provide a simple necessary and sufficient condition for a null hypothesis to satisfy the randomization hypothesis that avoids group-theoretic properties that typically accompany the theory of randomization tests.
We then apply our framework to analyze one-sample randomization tests. We show that for certain null hypotheses — such as the null of mean zero — there does not exist a randomization test that properly controls finite-sample error. These results are related to results concerning the non-existence of reasonable inference procedures of certain non-parametric hypotheses (Bahadur & Savage 1956, Lehmann & Loh 1990, Romano 2004). By focusing on randomization tests, we are able to develop conditions for the (non-)existence of finite-sample randomization tests for many null hypotheses and among different families of distributions. For instance, this finding affirms that existing applications of randomization tests are not inadvertently paying “too much” of a Type 1 error when using tests based on sign changes.
We also use our framework to inform researchers on how to look for and construct randomization tests for null hypotheses for which standard constructions are not appropriate. When the data is drawn from a finite support, we provide an explicit characterization of which null hypotheses admit a randomization test. When the data is instead drawn from a continuous support, we consider which null hypotheses admit a randomization test based on groups of linear actions. To the best of our knowledge, existing randomization tests are all based on such groups. We show that the set of one-sample null hypotheses that admit such a test are extremely limited: they only correspond to subsets of symmetric distributions and of normal distributions.
As part of identifying whether a null hypotheses admits a randomization test, we also show how to use our framework to explicitly construct randomization tests. Throughout the paper, we provide explicit examples — including a construction of a normality test that makes use of rotation symmetries — that may be of theoretic or applied interest.
These conclusions yield a clear positive implication that applied work that uses randomization tests should continue to do so in the current manner. Our largely negative results also affirm and motivate a line of work that considers the asymptotic properties of randomization tests (see, e.g., Romano (1989, 1990), DiCiccio & Romano (2017), Canay et al. (2017), Lei & Bickel (2021), Pouliot (2024), Bai et al. (2024)). These works develop asymptotic properties of specific tests (such as the above test of a mean of zero) or of randomization tests more generally. Finally, our results also suggest that one-sample randomization tests that are not testing for symmetry or normality must necessarily come from non-linear transformations.
The remainder of the paper is organized as follows. Section 2 reviews the randomization test procedure and develops our general framework. Section 3 applies our framework to the case of one-samples tests considering i.i.d. data supported on a finite set. Section 4 then considers the case where the data is supported on the real line. Section 5 concludes.
2 The randomization test
2.1 Construction of the randomization test
The construction follows Lehmann et al. (1986); see Zhang & Zhao (2023) for a recent formulation and discussion of the test. Let $X=(X_{1},\dots,X_{n})$ be data taking values in a sample space $\mathcal{X}$ with distribution $P\in\Omega$. Consider the problem of testing $H_{0}:P\in\Omega_{0}$, and suppose that there is a group $\mathbb{G}$ of measurable transformations $g:\mathcal{X}\to\mathcal{X}$ such that $gX\stackrel{{\scriptstyle d}}{{=}}X$ whenever $X\in P\in\Omega_{0}$. For ease of exposition, we develop the procedure when $\mathbb{G}$ is a finite group of size $M$.
Let $T:\mathcal{X}\to\mathbb{R}$ be any test statistic used to test the null. For every $x\in\mathcal{X}$, let $T^{(1)}(x;\mathbb{G})\leq\dots\leq T^{(M)}(x;\mathbb{G})$ be the ordered values of $T(gx)$ as $g$ varies in $\mathbb{G}$. Given nominal level $\alpha\in(0,1)$, let $k\equiv M-\lfloor M\alpha\rfloor,$ where $\lfloor M\alpha\rfloor$ denotes the largest integer less than or equal to $M\alpha$. Let $M^{+}(x)$ and $M^{0}(x)$ be the number of values $T^{(j)}(x)$ that are greater than $T^{(k)}(x)$ and equal to $T^{(k)}(x)$, respectively. Set
$a(x)=\frac{M\alpha-M^{+}(x)}{M^{0}(x)}.$
Define the randomization test as
$\phi(x;\mathbb{G})=\begin{cases}1&\text{if $T(x)>T^{(k)}(x;\mathbb{G})$}\\ a(x)&\text{if $T(x)=T^{(k)}(x;\mathbb{G})$}\\ 0&\text{if $T(x)<T^{(k)}(x;\mathbb{G})$}\end{cases},$ (1)
The following Theorem shows that the randomization test has nominal level $\alpha$. The proof we present is slightly different than the proof in Lehmann et al. (1986), in part due to our focus on the role of the group structure $\mathbb{G}$.
Theorem 1.
Let $\phi$ be the randomization test as described above, and suppose that for any $g\in\mathbb{G}$, $gX\stackrel{{\scriptstyle d}}{{=}}X$ whenever $X\in P\in\Omega_{0}$. Then $\operatorname{\mathbb{E}}_{P}[\phi(X;\mathbb{G})]=\alpha$ $\forall\;P\in\Omega_{0}$.
Proof of Theorem 1.
Consider any $X\sim P\in\Omega_{0}.$ Then
$\operatorname{\mathbb{E}}_{P}[\phi(X;\mathbb{G})]=\frac{1}{M}\operatorname{\mathbb{E}}_{P}\left[\sum_{g\in\mathbb{G}}\phi(X;\mathbb{G})\right]=\frac{1}{M}\operatorname{\mathbb{E}}_{P}\left[\sum_{g\in\mathbb{G}}\phi(gX;g\mathbb{G})\right],$
where the last equality used the form of $\phi$ as in (1) and that $X\stackrel{{\scriptstyle d}}{{=}}gX$ for all $g\in\mathbb{G}$. By construction, $\frac{1}{M}\sum_{g\in\mathbb{G}}\phi(gx;\mathbb{G})=\alpha$ for any $x\in\mathcal{X}$, so that $\frac{1}{M}\operatorname{\mathbb{E}}_{P}\left[\sum_{g\in\mathbb{G}}\phi(gX;\mathbb{G})\right]=\alpha.$ To complete the proof, it thus suffices to show that
$\operatorname{\mathbb{E}}_{P}\left[\sum_{g\in\mathbb{G}}\phi(gx;\mathbb{G})\right]=\operatorname{\mathbb{E}}_{P}\left[\sum_{g\in\mathbb{G}}\phi(gX;g\mathbb{G})\right].$ (2)
This follows from the group property that $g\mathbb{G}=\mathbb{G}$ for all $g\in\mathbb{G}$, and thus $\phi(gx;\mathbb{G})=\phi(gx;g\mathbb{G})$ for all $x$ and $g\in\mathbb{G}$, so that (2) holds. ∎
Remark 1.
Suppose that we did not require that the collection of invertible transformations $\mathbb{G}$ is a group. For (2) to hold generally, we require that $\mathbb{G}g=\mathbb{G}$ for all $g\in\mathbb{G}$. A basic property of groups is that if the set $\mathbb{G}$ is a collection of invertible transformations, then $\mathbb{G}g=\mathbb{G}$ for all $g\in\mathbb{G}$ if and only if $\mathbb{G}$ is a group (with composition as the operation). In this sense, the requirement that $\mathbb{G}$ is a group is both necessary and sufficient.
2.2 Framework
Having established the construction, we now develop a simple framework to examine which null hypotheses admit a group that is invariant to all distributions in the null.
Definition 1 (Randomization hypothesis).
We say $\Omega_{0}$ satisfies the randomization hypothesis if there exists a group $\mathbb{G}$ of measurable transformations $g:\mathcal{X}\to\mathcal{X}$ such that
$\Omega_{0}\subseteq\Omega_{\mathbb{G}}\subset\Omega,$ (3)
where $\Omega_{\mathbb{G}}\equiv\{P\in\Omega\;:\;gX\stackrel{{\scriptstyle d}}{{=}}X\;\forall\;g\in\mathbb{G}\ ,\ X\sim P\}.$
Remark 2.
The set inclusions in (3) enforce that $\Omega_{0}$ is invariant to $\mathbb{G}$ while also ensuring that $\mathbb{G}$ is not invariant to all distributions in $\Omega$. The latter condition ensures that the test has power by excluding trivial choices of groups (such as the identity group) that are invariant to all distributions in $\Omega$.
Our analysis will focus on characterizing which null hypotheses do and do not satisfy the randomization hypothesis. For those that do, we will be interested in determining a group that allows for an implementation of the randomization test procedure. The following result provides a simple tool to answer both of these questions in a way that avoids group-theoretic concepts.
Proposition 1.
$\Omega_{0}$ satisfies the randomization hypothesis if and only if there exists a bijective measurable function $f:\mathcal{X}\to\mathcal{X}$ such that
$\Omega_{0}\subseteq\Omega_{f}\subset\Omega,$ (4)
where $\Omega_{f}\equiv\big\{P\in\Omega\;:\;fX\stackrel{{\scriptstyle d}}{{=}}X,X\sim P\big\}.$ When this holds, a randomization test can be constructed by using the group generated by $f$ or any collection of functions that satisfy (4).
All omitted proofs can be found in Supplementary Appendix B. The proof of the forward direction is immediate and the proof of the reverse direction follows from showing that if $f$ is invariant to all distributions in $\Omega_{0}$, then so is $f^{-1}$. We can then show that $\Omega_{0}$ satisfies the randomization hypothesis using the group generated by $f$.
Proposition 1 shows that to study which null hypotheses satisfy the randomization hypothesis, it suffices to consider individual bijective functions. Given a bijective function that satisfies (4), we can then implement the randomization test procedure by taking the group generated by the function. We will later illustrate this approach with specific examples.
In the rest of the paper, we will apply these results to analyze one-sample tests. With some recycling of notation, we will accordingly consider $X=(X_{1},\dots,X_{n})$ with $X_{i}$ drawn i.i.d. from distribution $P\in\Omega$ supported on a (subset of) the real line. In this notation, it is useful to let $\Omega$ denote the set of distributions for $X_{i}$.
3 One-sample tests on a finite support
In this section, we suppose $X_{i}$ is drawn i.i.d. from $P$ supported on the finite set $\mathcal{A}\equiv\{\alpha_{1},\dots,\alpha_{K}\}\subset\mathbb{R}$. We can then represent $P$ by the associated function $p:\mathcal{A}\to[0,1]$ and represent $\Omega_{\text{discrete}}$ via
$\Omega_{\text{discrete}}\equiv\left\{p:\mathcal{A}\to[0,1]\text{ s.t. }\sum_{k=1}^{K}p(\alpha_{k})=1\right\}.$ (5)
By Proposition 1, a null hypothesis $\Omega_{0}\subset\Omega_{\text{discrete}}$ satisfies the randomization hypothesis if and only if there is some function $f:\mathcal{A}^{n}\to\mathcal{A}^{n}$ such that $\Omega_{0}\subseteq\Omega_{f}\subset\Omega_{\text{discrete}}$. On a finite support, we can explicitly write $\Omega_{f}$ as
$\Omega_{f}=\bigcap_{x\in\mathcal{A}}\left\{p\in\Omega_{\text{discrete}}:\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(f_{j}(x))\right\}.$ (6)
Each of the individual sets in this intersection corresponds to the invariances of simpler functions that equal $f$ for a single value and otherwise equal the identity. This provides an explicit characterization of Proposition 1 for one-sample tests over finite supports.
Proposition 2.
$\Omega_{0}\subset\Omega_{\text{discrete}}$ satisfies the randomization hypothesis if and only if there exists $x,y\in\mathcal{A}^{n}$ such that $y$ is not a permutation of $x$ and such that $\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(y_{j})$ for all $p\in\Omega_{0}$.
Proposition 2 is useful for two reasons. First, it can be used to identify bijective functions that satisfy the invariance. We can then construct randomization tests by taking the groups generated by the functions. To illustrate this, Example 1 of Supplementary Appendix A uses the result to construct the usual test of symmetry around zero and to construct a test of two points having equal mass.
Second, Proposition 2 highlights that the null hypotheses that satisfy the randomization hypothesis are only those that place equal mass on different sets of points (accounting for multiplicities). As such, it is natural to expect that null hypotheses that are not characterized by such features likely do not satisfy the randomization hypotheses. In what follows, we use Proposition 2 to show that null hypotheses about $k$-th moments (including the mean) and quantiles do not satisfy the randomization hypothesis. These results and our explicit proofs illustrate how a researcher can use our results to determine if a null hypothesis of interest does or does not satisfy the randomization hypothesis.
Proposition 3.
If $|\mathcal{A}|>4$, then for any positive integer $t$ and $\beta\in(\alpha^{t}_{1},\alpha^{t}_{K})$, $\Omega_{0}=\{P\in\Omega_{\text{discrete}}:\operatorname{\mathbb{E}}_{P}[X_{i}^{t}]=\beta\}$ does not satisfy the randomization hypothesis.
Proposition 4.
If $|\mathcal{A}|>2$, then for any $p\in(0,1)$ and $q\in(\alpha_{1},\alpha_{K})$, $\Omega_{0}=\{P\in\Omega_{\text{discrete}}:\operatorname{\mathbb{P}}[X_{i}\leq q]=p\}$ does not satisfy the randomization hypothesis.
The proofs of Propositions 3 and 4 follow by showing that whenever we have $x,y\in\mathcal{A}^{n}$ with the two vectors not being equal as sets, then we can construct an explicit distribution in the null that does not place equal mass on $x$ and $y$. The conclusions then follow by Proposition 2.
Remark 3.
The cardinality requirements in Propositions 3 and 4 are necessary to avoid edge cases. For example, when $|\mathcal{A}|=3,$ then for any set $\mathcal{A}=\{-L,0,L\}$, the null of mean zero is equivalent to the null of symmetric distributions, which satisfies the randomization hypothesis using any group of sign changes. The proof of Proposition 3 provides an edge case when $|\mathcal{A}|=4$. The proof of Proposition 4 reveals that the statement holds when $|\mathcal{A}|=2$ as long as we exclude $p=1/2$. When $|\mathcal{A}|=2$ and $p=1/2$, we can construct the null hypothesis by using any group of sign changes.
4 One-sample tests on a continuous support
In this section, we suppose $X_{i}$ is drawn i.i.d. from $P$ supported on the real line. In what follows, we first consider groups of linear transformations, and explicitly characterize the set of all null hypotheses that are invariant to any such group. We then consider arbitrary groups, and develop and apply a necessary condition for a null hypothesis to satisfy the randomization hypothesis.
4.1 Groups of linear transformations
A group of linear transformations $\mathbb{G}$ is one in which each group element $g$ corresponds to a matrix $A_{g}:\mathbb{R}^{n}\to\mathbb{R}^{n}$. The following result shows that if $\mathbb{G}$ is a group of linear transformations, then the set of distributions that is invariant to $\mathbb{G}$ is either a subset of the symmetric distributions with mean zero or of the Gaussian distributions.
Proposition 5.
Let $P\in\Omega_{\text{cts}}$, where $\Omega_{\text{cts}}$ is a subset of distributions supported on the real line. Let $\mathbb{G}$ be a group of linear transformations. Then exactly one of the following is true: 1. $\Omega_{\mathbb{G}}=\Omega_{\text{cts}}$, 2. $\Omega_{\mathbb{G}}=\{P\in\Omega_{\text{cts}}:\text{$P$ symmetric around $0$}\}$, 3. $\Omega_{\mathbb{G}}\subseteq\{N(\mu,\sigma^{2})\in\Omega_{\text{cts}}:\mu\in\mathbb{R},\sigma^{2}\in\mathbb{R}_{>0}\}$, or 4. $\Omega_{\mathbb{G}}=\emptyset$.
The proof follows by representing $gX$ via $AX$ for some matrix $A$. If the matrix is diagonal, then the only possible invariance is over the symmetric distributions. If the matrix is not diagonal, then $AX$ and $X$ are random vectors of i.i.d. random variables with at least one component of $AX$ featuring a non-trivial linear combination of multiple components of $X$. By the Darmois-Skitovich Theorem, this implies $X_{i}$ must be i.i.d. Gaussian.
Proposition 5 establishes that if we restrict to groups formed by linear maps, one-sample randomization tests only generalize permutation tests — which are invariant to all distributions in one-sample settings — in their ability to test if a distribution belongs to the symmetric distributions with mean zero or to a subset of the Gaussian distributions. The former can be implemented using the group of sign changes. Example 2 of Supplementary Appendix A shows how to construct randomization tests of the latter.
4.2 Arbitrary groups
We now allow the group to be arbitrary. The classic Bahadur & Savage (1956) result establishes the impossibility of testing null hypotheses that are too ‘dense’ within the space of all continuously distributed distributions. In the context of analyzing randomization tests, it turns out that a similar conclusion holds under a related local version of this denseness property. Let $\Omega_{\text{cts}}(\mathcal{B})$ be the set of distributions supported on $\mathcal{B}\subseteq\mathbb{R}$.
Definition 2 (Locally-dense).
A set $\Omega_{0}\subset\Omega_{\text{cts}}(\mathcal{B})$ is locally-dense if, roughly speaking, any density on a finite union of small intervals can be matched (up to scale) by some density in $\Omega_{0}$. Formally, it is locally-dense if there exists $L:\mathbb{N}\to\mathbb{R}_{+}$ such that for any intervals $I_{1},\dots,I_{m}$ with $|I_{i}|<L(m)$ and any density $p$ supported on these intervals, there exists $H\in\Omega_{0}$ with density $h$ and scaling $\alpha\in(0,1]$ satisfying $h(x)=\alpha p(x)$ on $\bigcup_{j=1}^{m}I_{j}$.
The following result shows that locally-dense null hypotheses do not satisfy the randomization hypothesis.
Proposition 6.
If $\Omega_{0}$ is locally-dense, then $\Omega_{0}$ does not satisfy the randomization hypothesis.
The proof of Proposition 6 requires developing additional notation to address technical considerations arising from working with continuous distributions. The proof makes use of Proposition 1 and shows that for any function $f$ such that $\Omega_{0}\subseteq\Omega_{f}$, local denseness implies that $\Omega_{f}=\Omega$. Intuitively, the local denseness property ensures that the null hypothesis is locally rich enough to represent any density at finitely many locations.
We now use Proposition 6 to show that the same negative results obtained in Section 3 also hold when considering continuous distributions or when considering continuous distributions over a fixed interval.
Proposition 7.
For any positive integer $t$ and $\beta\in\mathbb{R}$, $\Omega_{0}=\{P\in\Omega_{\text{cts}}(\mathbb{R}):\operatorname{\mathbb{E}}_{P}[X_{i}^{t}]=\beta\}$ is locally-dense and thus does not satisfy the randomization hypothesis.
To provide some intuition for this result, consider the null hypothesis that $\operatorname{\mathbb{E}}_{P}[X_{i}]=\beta$. To show this null hypothesis is locally-dense, consider any set of $m$ intervals with width bounded by some $\delta>0$ (so that $L(m)=\delta$). For any density $p$ on these intervals with mean $\gamma_{p}$ and $\alpha\in(0,1]$, a distribution $H$ on the reals with density $\alpha p(x)$ on the intervals will have mean $\alpha\gamma_{p}+(1-\alpha)\operatorname{\mathbb{E}}_{H}[X_{i}|\text{not in intervals}]$. Since there always are two additional intervals that respectively lie above and below $\beta$ that do not intersect with the $m$ intervals, we can judiciously choose the weighting $\alpha$ and the mass on these two intervals so that the resulting $H$ has mean $\beta$. This shows that the null hypothesis is locally-dense. The following results all make use of similar constructions.
Proposition 8.
For any $p\in(0,1)$ and $q\in\operatorname{\mathbb{R}}$, $\Omega_{0}=\{P\in\Omega_{\text{cts}}(\mathbb{R}):\operatorname{\mathbb{P}}[X_{i}\leq q]=p\}$ is locally-dense and thus does not satisfy the randomization hypothesis.
Remark 4.
To further illustrate the use of Proposition 6, we show in Supplementary Appendix B.3 that the statement in Proposition 7 holds if we replace $\Omega_{\text{cts}}(\mathbb{R})$ with $\Omega_{\text{cts}}([b_{0},b_{1}])$ for some interval $b_{0},b_{1}]$ and $\beta$ in the support. The same holds with Proposition 8 for any $q\in(b_{0},b_{1})$. In this appendix, we also show that the null hypothesis of the variance equaling some value is locally-dense and does not satisfy the randomization hypothesis.
5 Conclusion
Randomization tests yield exact finite-sample Type 1 error control when the null satisfies the randomization hypothesis. However, many randomization tests used in practice only provide such guarantees for nulls substantially stronger than the intended null. This paper developed a framework for characterizing which null hypotheses admit randomization tests with exact finite-sample Type 1 error control.
Applying our framework to one-sample problems revealed impossibilities and limited possibilities of uses of finite-sample randomization tests. Many standard null hypotheses, including the null of mean zero, cannot be tested using any randomization test while maintaining finite-sample validity. We further showed that linear group actions — the basis of all existing randomization tests — can only test subsets of symmetric or normal distributions in one-sample settings. These impossibility results confirm that practitioners using existing tests are not missing better alternatives. While our results establish that testing beyond symmetry or normality requires non-linear transformations, we also show that even such transformations cannot be used to test many common null hypotheses. The framework we developed not only delineates these boundaries but also provides constructive guidance for identifying when nulls admit randomization tests and explicitly building such tests when they exist. Overall, our findings validate the current practice of using standard randomization tests and further motivate focusing on the asymptotic validity of randomization tests.
SUPPLEMENTARY MATERIAL
Appendix A Examples of randomization tests
Example 1 (Construction of randomization tests over discrete support).
  1. 1.
    For $L>0$, let $\mathcal{A}\equiv\{-L,\dots,0,\dots,L\}$ and let $\Omega_{0}$ be the set of distributions that are symmetric around 0. Consider any $x,y$ with $x_{j}\neq 0$ for some $j$ and y=xy=-x. We then have that $c(x_{j};x)>c(x_{j};y)$ and that for all $p\in\Omega_{0}$, $\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(-x_{j})=\prod_{j=1}^{n}p(y_{j})$. By Proposition 2 $\Omega_{\text{discrete}}$ satisfies the randomization test. We can then construct a randomization test for $\Omega_{0}$ by taking the group arising from any collection of functions that send some values of $x\in\mathcal{A}^{n}$ to x-x and otherwise act as the identity function.
  2. 2.
    For any $\mathcal{A}$, let $\Omega_{0}$ be the set of distributions where $p(\alpha_{\ell})=p(\alpha_{\ell^{\prime}})$ for some $\alpha_{\ell}\neq\alpha_{\ell^{\prime}}$. Consider any $x,y$ with $x$ containing $\alpha_{\ell}$ but not $\alpha_{\ell^{\prime}}$ and $y_{j}=\alpha_{\ell^{\prime}}\mathbb{1}[x_{j}=\alpha_{\ell}]+\alpha_{\ell}\mathbb{1}[x_{j}=\alpha_{\ell^{\prime}}]+x_{j}\mathbb{1}[x_{j}\neq\alpha_{\ell},x_{j}\neq\alpha_{\ell^{\prime}}]$. We then have that $c(\alpha_{\ell};x)>c(\alpha_{\ell};y)$ and for all $p\in\Omega_{0}$, $\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(y_{j})$. By Proposition 2 $\Omega_{\text{discrete}}$ satisfies the randomization test. We can then construct a randomization test for $\Omega_{0}$ by taking the group arising from any collection of functions that acts as the identity except for swapping the values of $\alpha_{\ell}$ and $\alpha_{\ell^{\prime}}$.
Example 2 (Randomization tests for Gaussian distributions).
  1. 1.
    Let $\mathbb{G}$ be the orthogonal group of degree $n$. By Maxwell’s Theorem, the set of invariances given $\mathbb{G}$ is $\Omega_{\mathbb{G}}=\{N(0,\sigma^{2}):\sigma\in\mathbb{R}\}.$ Thus, this group can be used to construct a randomization test of a distribution being Gaussian with mean zero.
  2. 2.
    Let $\mathbb{G}$ be the set of transformations generated by $3\times 3$ block permutations of the matrix
    $\begin{pmatrix}\frac{2}{3}&-\frac{1}{3}&\frac{2}{3}\\ \frac{2}{3}&\frac{2}{3}&-\frac{1}{3}\\ -\frac{1}{3}&\frac{2}{3}&\frac{2}{3}\\ \end{pmatrix}$
    In the below proof, we show that $\mathbb{G}$ is a group and that $\Omega_{\mathbb{G}}=\{N(\mu,\sigma^{2}):(\mu,\sigma)\in\mathbb{R}^{2}\}.$ Thus, this group can be used to construct a randomization test of a distribution being Gaussian.
Proof of claims in Example 2.
Let $X_{1},X_{2},X_{3}$ be i.i.d. draws from $N(\mu,\sigma^{2})$ for some $\mu,\sigma\in\operatorname{\mathbb{R}}$. Let $X=(X_{1},X_{2},X_{3})$ and $V=(\mu,\mu,\mu)$. Because
$A\equiv\begin{pmatrix}\frac{2}{3}&-\frac{1}{3}&\frac{2}{3}\\ \frac{2}{3}&\frac{2}{3}&-\frac{1}{3}\\ -\frac{1}{3}&\frac{2}{3}&\frac{2}{3}\\ \end{pmatrix}$
is orthogonal, we know from Maxwell’s Theorem,
$A(X-V)\stackrel{{\scriptstyle d}}{{=}}X-V.$
Since $AV=V$, we have that
$AX\stackrel{{\scriptstyle d}}{{=}}X.$
Inductively, $A^{n}X\stackrel{{\scriptstyle d}}{{=}}X$ for all $n\in\mathbb{Z}$. Thus, $\{N(\mu,\sigma^{2}):\mu,\sigma\in\operatorname{\mathbb{R}}\}\subset\Omega_{\mathbb{G}}$. By proposition 5, $\{N(\mu,\sigma^{2}):\mu,\sigma\in\operatorname{\mathbb{R}}\}=\Omega_{\mathbb{G}}$. ∎
Appendix B Proofs
Appendix B.1 provides the proofs of all propositions except for the proof of Proposition 6. The proof of Proposition 6 requires additional notation and definitions and is developed in Appendix B.2. Lastly, Appendix B.3 provides statements and proofs of the claims made in Remark 4.
B.1 All proofs excluding the proof of Proposition 6
Proof of Proposition 1.
The forward direction follows from noting that $\Omega_{\mathbb{G}}=\cap_{g\in\mathbb{G}}\Omega_{g}$, and since (3) holds, there must be at least some $f\in\mathbb{G}$ such that (4) holds. Group elements have inverses and thus $f$ is bijective.
For the reverse direction, consider a bijective measurable $f$ satisfying (4). Note that the set generated by $f$, denoted $\langle f\rangle\equiv\{f^{K}:K\in\mathbb{Z}\}$, is a group with composition as the operation, with $f^{-K}={f^{-1}}^{K}$ where $f^{-1}$ denotes the inverse of $f$. Then $\Omega_{\langle f\rangle}\subseteq\Omega_{f}\subset\Omega$.
It remains to show that $\Omega_{0}\subseteq\Omega_{\langle f\rangle}$, which is equivalent to showing that $\operatorname{\mathbb{P}}[f^{K}X\in A]=\operatorname{\mathbb{P}}[X\in A]$ for all $X\sim P\in\Omega_{0}$, $K\in\mathbb{Z}$, and measurable $A$. When $K=0$, the proof follows trivially since the identity is invariant to all distributions. For any integer $K>0$,
$\operatorname{\mathbb{P}}[f^{K}X\in A]=\operatorname{\mathbb{P}}[fX\in f^{K-1}A]=\operatorname{\mathbb{P}}[X\in f^{K-1}A]=\operatorname{\mathbb{P}}[f^{K-1}X\in A],$ (7)
where we used that $f$ is invertible and also used that $f$ is measurable so that $f^{K-1}A$ is measurable and thus, since $\Omega_{0}\subseteq\Omega_{f}$, we have $\operatorname{\mathbb{P}}[fX\in f^{K-1}A]=\operatorname{\mathbb{P}}[X\in f^{K-1}A]$. Repeating this step, we obtain $\operatorname{\mathbb{P}}[f^{K}X\in A]=\operatorname{\mathbb{P}}[X\in A]$. Lastly, note that
$\operatorname{\mathbb{P}}[f^{-1}X\in A]=\operatorname{\mathbb{P}}[X\in fA]=\operatorname{\mathbb{P}}[fX\in fA]=\operatorname{\mathbb{P}}[X\in A],$ (8)
where the second equality used that $\Omega_{0}\subseteq\Omega_{f}$ and that $fA$ is measurable so that $\operatorname{\mathbb{P}}[X\in fA]=\operatorname{\mathbb{P}}[fX\in fA]$. Then for $K>0$, arguments as above imply
$\operatorname{\mathbb{P}}[f^{-K}X\in A]=\operatorname{\mathbb{P}}[f^{-1}X\in f^{-(K-1)}A]=\operatorname{\mathbb{P}}[X\in f^{-(K-1)A}]=\operatorname{\mathbb{P}}[f^{-(K-1)}X\in A].$ (9)
Repeating this step, we obtain $\operatorname{\mathbb{P}}[f^{-K}X\in A]=\operatorname{\mathbb{P}}[X\in A]$, as required. ∎
Proof of Proposition 2.
By Proposition 1, $\Omega_{0}$ satisfies the randomization hypothesis if and only if there exists a bijective $f$ such that $\Omega_{0}\subseteq\Omega_{f}\subset\Omega_{\text{discrete}}$, where we drop measurability since all functions are measurable over discrete sets. Expanding the form of $\Omega_{f}$ using (6), we have that
$\Omega_{f}=\bigcap_{x\in\mathcal{A}}\left\{p\in\Omega_{\text{discrete}}:\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(y_{j})\ \text{s.t.}\ y=f(x)\right\}.$ (10)
Then any $f$ such that $\Omega_{0}\subseteq\Omega_{f}\subset\Omega_{\text{discrete}}$ must have some $x^{\star}\in\mathcal{A}$ and $y^{\star}=f(x^{\star})$ such that
$\Omega_{0}\subseteq\underbrace{\left\{p\in\Omega_{\text{discrete}}:\prod_{j=1}^{n}p(x^{\star}_{j})=\prod_{j=1}^{n}p(y^{\star}_{j})\right\}}_{\equiv\Omega^{\star}}\subset\Omega_{\text{discrete}}.$ (11)
If $x^{\star}$ and $y^{\star}$ are permutations of each other so that $y^{\star}=(x^{\star}_{\sigma(1)},\dots,x^{\star}_{\sigma(n)})$ for some permutation $\sigma\in S_{n}$, then $\prod_{i=1}^{n}p(x^{\star}_{j})=\prod_{j=1}^{n}p(y^{\star}_{j})$ for all $p\in\Omega_{\text{discrete}}$, so $\Omega^{\star}=\Omega_{\text{discrete}}$. It thus must be that these two vectors are not permutations of each other. The proof is completed by showing that when this is not the case, there is some $p\in\Omega_{\text{discrete}}$ that does not belong to $\Omega^{\star}$. We can construct such a $p$ by noting that there is at least some element $\alpha\in\mathcal{A}$ that appears more times in $x^{\star}$ than in $y^{\star}$, and we can then define $p$ as placing mass $\epsilon_{0}$ on all values except $\alpha$ and mass $\epsilon_{1}\neq\epsilon_{0}$ on $\alpha$, so that $\prod_{j=1}^{n}p(x^{\star}_{j})\neq\prod_{j=1}^{n}p(y^{\star}_{j})$, as required. ∎
Proof of Proposition 3.
Define $c(\alpha;x)$ as the number of times the value $\alpha$ shows up in the vector $x$. To show that $\Omega_{0}=\{p\in\Omega_{\text{discrete}}:\operatorname{\mathbb{E}}_{p}[X_{i}^{t}]=\beta\}$ does not satisfy the randomization hypothesis, it suffices to show that for all $x,y\in\mathcal{A}^{n}$ with $c(\alpha;x)>c(\alpha;y)$ for some $\alpha\in\mathcal{A}$, there exists some $p\in\Omega_{0}$ such that $\prod_{j=1}^{n}p(x_{j})\neq\prod_{j=1}^{n}p(y_{j})$.
Let $x,y\in\mathcal{A}^{n}$ be arbitrary such that $c(\alpha;x)>c(\alpha;y)$ for some $\alpha\in\mathcal{A}$.
First, consider the case where $t$ is odd. For simplicity, let us assume that $\alpha_{1}<\alpha_{2}<\cdots<\alpha_{K}$. For $x,y\in\mathcal{A}^{n}$, we define $\beta_{i}(x,y)=c(\alpha_{i};x)-c(\alpha_{i};y)$ and notice that $\sum_{i=1}^{K}\beta_{i}(x,y)=0$. Let’s consider the following two cases:
  1. 1.
    There exists some $1\leq i\leq K-1$ such that $\alpha_{i}^{t}<\beta<\alpha^{t}_{i+1}$. Consider the following subcases:
    1. (a)
      There exists some $i_{1}\leq i$ and $i_{2}\geq i+1$ such that $|\beta_{i_{1}}(x,y)|\neq|\beta_{i_{2}}(x,y)|$. Let $p(\alpha_{j})=\epsilon$ for all $j\neq i_{1},i_{2}$, $p(\alpha_{i_{1}})=q$, and $p(\alpha_{i_{2}})=1-(K-2)\epsilon-q$. By letting $\epsilon$ to be arbitrarily small, we can ensure $\prod_{j=1}^{K}p(x_{j})\neq\prod_{j=1}^{K}p(y_{j})$ as one of the products will have more $\epsilon$ factors. We then adjust $q$ such that $\sum_{j=1}^{K}\alpha_{j}^{t}p(\alpha_{j})=\beta$.
    2. (b)
      For all $1\leq j\leq K$, $|\beta_{j}(x,y)|=C>0$ for some positive integer $C$. Because $|\mathcal{A}|>2$, we can choose $\alpha_{i_{1}}$ and $\alpha_{i_{2}}$ such that $|\alpha^{t}_{i_{1}}-\beta|\neq|\alpha^{t}_{i_{2}}-\beta|$. Then let $p(\alpha_{j})=\epsilon$ for $j\neq i_{1},i_{2}$, $p(\alpha_{i_{1}})=q$, and $p(\alpha_{i_{2}})=1-(K-2)\epsilon-q$. The argument above still applies because in order for $\sum_{j=1}^{K}\alpha_{j}^{t}p(\alpha_{j})=\beta$ to hold, we need $q\approx\frac{\beta-\alpha_{i_{2}}^{t}}{\alpha_{i_{1}}^{t}-\alpha_{i_{2}}^{t}}.$ Since $|\alpha^{t}_{i_{1}}-\beta|\neq|\alpha^{t}_{i_{2}}-\beta|$, $1-(K-2)\epsilon-q\neq q$. Although $\prod_{j=1}^{K}p(x_{j})$ and $\prod_{j=1}^{K}p(y_{j})$ may have the same number of $\epsilon$ factors, $1-(K-2)\epsilon-q\neq q$ prevents the two product to be equal.
  2. 2.
    There exists some $2\leq i\leq K-1$ such that $\beta=\alpha_{i}^{t}$. If $\beta_{i}(x,y)\neq 0$, we can just let $p(\alpha_{i})=1-\sum_{j=1,j\neq i}^{K}\epsilon_{j}$ and $p(\alpha_{j})=\epsilon_{j}$ for $j\neq i$. By letting $\epsilon_{j}$ to be arbitrarily small, we can still ensure that $\sum_{k=1}^{K}\alpha_{k}^{t}p(\alpha_{k})=\beta$ and $\prod_{k=1}^{K}p(x_{k})\neq\prod_{k=1}^{K}p(y_{k})$. For $x,y\in\mathcal{A}^{n}$ such that $\beta_{i}(x,y)=0$, we just repeat the argument for case 1 without choosing $i_{1}$ or $i_{2}$ to be $i$ and here we need $|\mathcal{A}|>3$.
When $t$ is even, we can still relabel $\alpha_{i}$ such that $\alpha_{1}^{t}\leq\alpha_{2}^{t}\leq\cdots\leq\alpha_{K}^{t}.$ With the requirement that $|\mathcal{A}|>4$, we can easily generalizes the argument for odd moments except for the problematic case where $\beta=\alpha_{i}^{t}$, $\beta_{i}(x,y)=0$, $|\beta_{j}(x,y)|=C>0$ for all $1\leq j\leq K$ and $j\neq i$, and $|\alpha_{j}^{t}-\beta|=D>0$ for all $j\neq i$. In this case, $|\{\alpha_{i}^{t}|1\leq i\leq K\}|=3$. Since $|\mathcal{A}|>4$ and $\sum_{j=1}^{K}\beta_{j}(x,y)=0$, we can only have $\alpha_{1}^{t}=\alpha_{2}^{t}$ and $\alpha_{K-1}^{t}=\alpha_{K}^{t}$. Then we can let $p(\alpha_{1})=p(\alpha_{2})=1/4-1/4\epsilon$, $p(\alpha_{K})=1/2-1/2\epsilon-\delta$ and $p(\alpha_{K-1})=\delta$ and $p(\alpha_{i})=\epsilon$. By letting $\epsilon$ and $\delta$ to be arbitrarily small, we can repeat the same argument as above.
When $|\mathcal{A}|=4$, the set $\mathcal{A}=\{-1,0,1,\sqrt{2}\}$ with the null of the second moment equaling one is equivalent to the null that 0 and $\sqrt{2}$ have equal mass. A randomization test can be constructed for this latter null by permuting these two numbers (see Example 1 of Appendix A).
Proof of Proposition 4.
Let $x,y\in\mathcal{A}^{n}$ be arbitrary such that $c(\alpha;x)>c(\alpha;y)$ for some $\alpha\in\mathcal{A}$ and let $\beta_{i}(x,y)=c(\alpha_{i};x)-c(\alpha_{i};y)$. We know that there exists some $1\leq i\leq K-1$ such that $\alpha_{i}\leq q<\alpha_{i+1}$. One of the follow two cases must occur:
  1. 1.
    There exists some $i_{1}\leq i$ and $i_{2}>i$ such that $|\beta_{i_{1}}(x,y)|\neq|\beta_{i_{2}}(x,y)|$. Then we let $p(\alpha_{j})=\epsilon$ for $j\neq{i_{1}},{i_{2}}$, $p(\alpha_{i_{1}})=p-(i-1)\epsilon$, and $p(\alpha_{i_{2}})=1-(K-i-1)\epsilon-p$. One of the $\prod_{j=1}^{K}p(x_{j})$ and $\prod_{j=1}^{K}p(y_{j})$ will have more $\epsilon$ factors and hence are not the same.
  2. 2.
    For all $1\leq j\leq K$, $|\beta_{j}(x,y)|=C>0$ for some constant $C$. Because $|\mathcal{A}|>2$, without loss of generality, we can assume that there exists $i_{1},i_{2},i_{3}$ such that $i_{1}<i_{2}\leq i<i_{3}$. Then we let $p(\alpha_{j})=\epsilon$ for $j\neq{i_{1}},{i_{2}},{i_{3}}$, $p(\alpha_{i_{1}})=\frac{1}{2}p$, $p(\alpha_{i_{2}})=\frac{1}{2}p-(i-2)\epsilon$, and $p(\alpha_{i_{3}})=1-(K-i-1)\epsilon-p$. The argument above still applies because one of the $\prod_{j=1}^{K}p(x_{j})$ and $\prod_{j=1}^{K}p(y_{j})$ will have more $\epsilon$ factors.
Proof of Proposition 5.
Since $\mathbb{G}$ is a group of linear maps, any element of the group is a bijective and thus invertible $n\times n$ matrix $A=\{a_{ij}\}_{ij}$. Suppose there exists $i,i^{\prime}$ and some $j^{\prime}$ such that $a_{ij^{\prime}},a_{i^{\prime}j^{\prime}}\neq 0$. If $\Omega_{A}$ is not empty, then
$X_{i}=\sum_{j}a_{ij}X_{j}$ (12)
$X_{i^{\prime}}=\sum_{j}a_{i^{\prime}j}X_{j},$ (13)
where $\{X_{j}\}_{j}$ are independent and $X_{i},X_{i^{\prime}}$ are also independent (otherwise $AX\not\stackrel{{\scriptstyle d}}{{=}}X$). We can thus apply the Darmois-Skitovich Theorem (for a reference, see Chapter 3 of Kagan et al. 1973), which implies that since $a_{ij^{\prime}}a_{i^{\prime}j^{\prime}}\neq 0$, $X_{j^{\prime}}$ is Gaussian. It follows that $X_{1},\dots,X_{n}$ are i.i.d. Gaussian, so that $\Omega_{A}$ is a subset of Gaussian distributions.
Alternatively, suppose there does not exist $i,i^{\prime}$ and some $j^{\prime}$ such that $a_{ij^{\prime}},a_{i^{\prime}j^{\prime}}\neq 0$. Then for any $j$, $\{a_{1j},\dots,a_{nj}\}$ is zero for all entries except for some index $i_{j}$. Since $A$ is invertible, for each row, at least one entry must be non-zero. Since each column has exactly one non-zero entry and since each row must have one non-zero entry, it must be that
$\begin{bmatrix}d_{1}&0&\dots&0\\ 0&d_{2}&0&\vdots\\ \vdots&0&\ddots&0\\ 0&\dots&0&d_{n}\end{bmatrix}\sigma_{n}$ (14)
where $d_{i}\neq 0$ and $\sigma_{n}$ is a permutation of the indices of $(x_{1},\dots,x_{n})$. Then, given $i$, $X_{i}=d_{j}X_{j}$ for some $j$. Then $\operatorname{Var}[X_{i}]=d_{j}^{2}\operatorname{Var}[X_{i}]$, and thus for $\Omega_{A}$ to be non-empty, it must be that $d_{j}\in\{-1,1\}$. If $d_{j}=-1$, then for $X_{i}$ with underlying distribution $P$ to lie in $\Omega_{A}$, it must be that $\operatorname{\mathbb{E}}[X_{i}]=-\operatorname{\mathbb{E}}[X_{i}]$ so that $\operatorname{\mathbb{E}}[X_{i}]=0$ and that $X_{i}=-X_{i}$, which is satisfied if and only if $P$ is symmetric around zero. Thus $\Omega_{A}$ equals the distributions that are symmetric around zero. Finally, if $d_{j}=1$, then $X_{i}=X_{i}$ is satisfied by any distribution, so that $\Omega_{A}=\Omega$. Since $\mathbb{G}$ is a collection of linear invertible matrix, $\Omega_{\mathbb{G}}$ is an intersection of these possible sets of distributions, as required. ∎
Proof of Proposition 7.
Let $X$ be a continuous random variable with probability density function $p$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]$. We will show that for any $\beta\in\operatorname{\mathbb{R}}$ we can find an $\alpha\in(0,1]$ and a probability density function $g$ on the real line with $(2k+1)$-th moment equal to $\beta$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$. Let $h$ be a probability density function supported on $\operatorname{\mathbb{R}}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=ap(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then
$\operatorname{\mathbb{E}}[X_{a}^{2k+1}]=a\operatorname{\mathbb{E}}[X^{2k+1}]+(1-a)\operatorname{\mathbb{E}}[Y^{2k+1}].$
Let $P$ be a positive number such that $P^{2k+1}>\beta$ and $Q$ be a negative number such that $Q^{2k+1}<\beta$. Let
$h(x)=\begin{cases}\mathbb{1}_{[P,P+1]}&\operatorname{\mathbb{E}}[X^{2k+1}]\leq\beta\\ \mathbb{1}_{[Q-1,Q]}&\operatorname{\mathbb{E}}[X^{2k+1}]>\beta\end{cases}$ (15)
Then,
$\operatorname{\mathbb{E}}[Y^{2k+1}]=\begin{cases}\frac{1}{2k+2}\sum_{i=0}^{2k+1}P^{i}(P+1)^{{2k+1}-i}\geq P^{2k+1}>\beta&\operatorname{\mathbb{E}}[X^{2k+1}]\leq\beta\\ \frac{1}{2k+2}\sum_{i=0}^{2k+1}(Q-1)^{i}Q^{{2k+1}-i}\leq Q^{2k+1}<\beta&\operatorname{\mathbb{E}}[X^{2k+1}]>\beta.\end{cases}$ (16)
In both cases, we can find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X_{\alpha}^{2k+1}]=\beta$.
Thus, for any $\beta\in\operatorname{\mathbb{R}}$ and $k\in\mathbb{Z}_{\geq 0}$, $\{P\in\Omega_{\text{cts}}:\operatorname{\mathbb{E}}_{P}[X_{i}^{2k+1}]=\beta\}$ is locally-dense.
We now consider the $2k$-th moment, and thus require that $\beta>0$ (else the null hypothesis is empty). Let $X$ be a continuous random variable with probability density function $p$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]$ where $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\beta^{\frac{1}{2k}}}{m+1}$. We can find an $\alpha\in(0,1]$ and a probability density function $g$ on the real line with $2k$-th moment equal to $\beta$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$. Notice that the choice of the upper bound for the interval length plays a crucial role in the definition of the locally-dense. If we remove that condition from the definition, then we can choose a probability density function $p(x)$ to be supported on $\operatorname{\mathbb{R}}\setminus[-2\beta^{\frac{1}{2k}},2\beta^{\frac{1}{2k}}]$. In this case, whatever probability density function $g$ and $\alpha\in(0,1]$ we choose such that $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$, the $2k$-th moment of the random variable with density $g$ is strictly greater than $\beta$.
Again, let $h$ be a probability density function supported on $\operatorname{\mathbb{R}}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=ap(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then
$\operatorname{\mathbb{E}}[X_{a}^{2k}]=a\operatorname{\mathbb{E}}[X^{2k}]+(1-a)\operatorname{\mathbb{E}}[Y^{2k}].$
If $\operatorname{\mathbb{E}}[X^{2k}]\leq\beta$, let $P$ be a positive number such that $P^{2k}>\beta$ and $P,P+1]\subset(\operatorname{\mathbb{R}}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}])$ and let
$h(x)=\mathbb{1}_{[P,P+1]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k}]\geq P^{2k}>\beta,$
we can easily find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k}_{\alpha}]=\beta$.
On the other hand, if $\operatorname{\mathbb{E}}[X^{2k}]>\beta$, given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\beta^{\frac{1}{2k}}}{m+1}$, we can find an interval $x,y]\subset[0,\beta^{\frac{1}{2k}}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Let
$h(x)=\frac{1}{y-x}\mathbb{1}_{[x,y]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k}]\leq y^{2k}\leq\beta.$
We can then find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k}_{\alpha}]=\beta$.
Thus, for any $\beta\in\operatorname{\mathbb{R}}_{>0}$ and $k\in\mathbb{Z}_{\geq 1}$, $\{P\in\Omega_{\text{cts}}:\operatorname{\mathbb{E}}_{P}[X_{i}^{2k}]=\beta\}$ is locally-dense. ∎
Proof of Proposition 8.
Let $X$ be a continuous random variable with probability density function $f$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]$. We will show that for any $p\in(0,1)$ and $q\in\operatorname{\mathbb{R}}$, we can find an $\alpha\in(0,1]$ and a probability density function $g$ on the real line with the $p$-th quantile equal to $q$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha f(x)$.
Let $h$ be a probability density function supported on $\operatorname{\mathbb{R}}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=af(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then,
$\operatorname{\mathbb{P}}[X_{a}\leq q]=a\operatorname{\mathbb{P}}[X\leq q]+(1-a)\operatorname{\mathbb{P}}[Y\leq q].$
There exists a large integer $N$ such that $x_{1},y_{1}],\cdots,[x_{m},y_{m}],q\subset[-N+1,N-1]$. Choose $\alpha>0$ small enough such that $\alpha\operatorname{\mathbb{P}}[X\leq q]<p$. In order for $\operatorname{\mathbb{P}}[X_{\alpha}\leq q]=p$, we need $\operatorname{\mathbb{P}}[Y\leq q]=\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}.$ This can be achieved by letting $h(x)=\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}\mathbb{1}_{[-N-1,-N]}+(1-\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha})\mathbb{1}_{[N,N+1]}$.
B.2 Proof of Proposition 6
We prove Proposition 6 by making use of Proposition 1 and showing that if $\Omega_{0}$ is locally-dense, then there is no function that is invariant to all distributions in $\Omega_{0}$ that is also not invariant to all distributions in $\Omega$. To do so, we partition the space of functions into three classes of functions. The first are ‘approximate permutation’ maps that behave like permutation maps. We will show that these functions are invariant to all continuous distributions, so cannot be used to construct a randomization test. The other two classes of functions will satisfy properties that allow us to construct, for any such function, a distribution in $\Omega_{0}$ that is not invariant to the function, which would yield the result.
B.2.1 Definitions
We begin by defining some notations and concepts. For a subset $A\subseteq\operatorname{\mathbb{R}}^{n}$, we let $m(A)$ denote the Lebesgue measure of the set. We let $\sigma_{1},\cdots,\sigma_{n!}$ be the distinct $n!$ permutations of coordinates. Let $A_{i}=\{\sigma_{i}(x):x\in A\}$ be the image of $A$ under the permutation map $\sigma_{i}$. Let $\bar{A}=\cup_{i=1}^{n!}A_{i}$ be the union of all the images of $A$ under permutation maps.
Definition 3.
We call a bijective function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$ mass-preserving if for any measurable set $A\subset\operatorname{\mathbb{R}}^{n}$, $m(A)=0$ if and only if $m(f(A))=0$.
Definition 4.
A bijective function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$ is orbit-preserving if for any measurable set $A\subset\operatorname{\mathbb{R}}^{n}$, $m(f(A)\setminus\bar{A})=0$.
Definition 5.
A bijective function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$ is box-orbit-preserving if for any box $A=\prod_{i=1}^{n}[a_{i},b_{i}]\subset\operatorname{\mathbb{R}}^{n}$, $m(f(A)\setminus\bar{A})=0$.
Definition 6.
We call a bijective function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$ an approximate permutation map if $f$ is box-orbit-preserving and mass-preserving.
Notice that both orbit-preserving and box-orbit-preserving properties imply that for any measurable set $A\subset\operatorname{\mathbb{R}}^{n}$, the image of $A$ under the approximate permutation maps are “contained” in the image of $A$ under the actual permutation maps, not the other way around. However, it suffices to require only box-orbit-preserving and mass-preserving properties to define approximate permutation maps because these properties apply to all the measurable sets $A$ and this turns out to be exactly as strong as we need.
B.2.2 Proof strategy
A function is either an approximate permutation map, or not mass-preserving, or mass-preserving but not box-orbit-preserving. The proof of Proposition 6 will follow by showing that functions that are approximate permutation maps are invariant to all distributions, while functions in the latter two classes must always fail to be invariant to some distribution in $\Omega_{0}$ when it is locally-dense. We consider each space of functions separately in the following series of lemmas, which will immediately obtain the main result. The proofs of the Lemmas are provided in Appendix B.2.3 to not disrupt the flow of the arguments here.
Lemma 1.
A box-orbit-preserving map is orbit-preserving.
Lemma 2.
An approximate permutation map is invariant to all i.i.d. continuous distributions.
We next consider functions that are not approximate permutation maps. These functions are either not mass-preserving or mass-preserving but not box-orbit-preserving. For such functions, we will essentially find several small intervals on which the map $f$ acts very differently from the permutation maps. On those intervals, we can define distributions in $\Omega_{0}$ that are not preserved by $f$. Outside those intervals, we will define distributions based on the locally-dense property. Consequently, when we combine these distributions, we will have a $f$-specific distribution that is not invariant to $f$ and belongs to $\Omega_{0}$.
Lemma 3.
For any function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$ that is not mass-preserving, there exists a random vector $X=(X_{1},\dots,X_{n})$, $X_{i}$ drawn i.i.d. from distribution $P\in\Omega_{0}$ where $\Omega_{0}$ is locally-dense such that $X$ is not invariant under $f$.
Lemma 4.
For any mass-preserving and not box-orbit-preserving function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$, there exists a random vector $X=(X_{1},\dots,X_{n})$, $X_{i}$ drawn i.i.d. from distribution $P\in\Omega_{0}$ where $\Omega_{0}$ is locally-dense such that $X$ is not invariant under $f$.
The proof of Proposition 6 follows immediately from the above lemmas.
Proof of Proposition 6.
By definition, a function is either an approximate permutation map, not mass-preserving, or mass-preserving but not box-orbit-preserving. If it is an approximate permutation map, Lemma 1 shows it is invariant to all distributions. If it is not, Lemmas 2-4 show it is not invariant to all distributions in $\Omega_{0}$ when it is locally-dense. The result then follows by Proposition 1. ∎
B.2.3 Proof of lemmas
Proof of Lemma 1.
Assume for contradiction that $f$ is box-orbit-preserving and not orbit-preserving. Then there exists a measurable set $A\subset\operatorname{\mathbb{R}}^{n}$ such that $m(f(A)\setminus\bar{A})=c>0$. Let $B=f(A)\setminus\bar{A}$ and note that $B\cap\overline{f^{-1}(B)}=\emptyset$. Since $f$ is mass-preserving, $m(f^{-1}(B))>0$. We can find a finite number of pairwise disjoint boxes $D_{1},\cdots,D_{k}$ where $D_{j}=\prod_{i=1}^{n}[a^{j}_{i},b^{j}_{i}]$ such that $f^{-1}(B)\subset\cup_{j=1}^{k}D_{j}$ and $m(\cup_{j=1}^{k}D\setminus f^{-1}(B))<\frac{c}{2n!}$. Let $D=\cup_{j=1}^{k}D_{j}$.
$m(f(D)\setminus\bar{D})\geq m(B\setminus\bar{D})=m(B\setminus(\overline{D\setminus f^{-1}(B)}))\geq m(B)-n!m(D\setminus f^{-1}(B))\geq\frac{c}{2}>0.$
Thus, there exists at least one $D_{j}$ such that $m(f(D_{j})\setminus\bar{D_{j}})>0$. ∎
Proof of Lemma 2.
First, for any measurable set $A\subset\operatorname{\mathbb{R}}^{n}$, we can find a disjoint partition of $f(A)=\bigsqcup_{i=1}^{n!}B_{i}\bigsqcup C$ such that $B_{i}\subset\sigma_{i}(A)$, $C\subset\bar{A}^{c}$. By Lemma 1, we know that an approximate permutation map is also orbit-preserving. Thus, $m(f(A)\setminus\bar{A})=0$ and $m(C)=0$. Let $A_{i}=f^{-1}(B_{i}).$ We would like to show that $\sigma_{i}(A_{i})=B_{i}$ almost everywhere, i.e., $m(\sigma_{i}(A_{i})\triangle B_{i})=0$, where $\sigma_{i}(A_{i})\triangle B_{i}$ denotes the symmetric difference between two sets $A_{i}$ and $B_{i}$. Then, it follows
$\begin{split}\operatorname{\mathbb{P}}[X\in f(A)]=\sum_{i=1}^{n!}\operatorname{\mathbb{P}}[X\in B_{i}]=\sum_{i=1}^{n!}\operatorname{\mathbb{P}}[X\in\sigma_{i}(A_{i})]=\sum_{i=1}^{n!}\operatorname{\mathbb{P}}[X\in A_{i}]=\operatorname{\mathbb{P}}[X\in A].\end{split}$
Now to show that $m(\sigma_{i}(A_{i})\triangle B_{i})=0$, we need to show $m(\sigma_{i}(A_{i})\setminus B_{i})=0$ and $m(B_{i}\setminus\sigma_{i}(A_{i}))=0$. The latter one is trivial as $m(B_{i}\setminus\sigma_{i}(A_{i}))=m(f(A_{i})\setminus\sigma_{i}(A_{i}))=0$ follows from the condition given. It is only left to show that $m(\sigma_{i}(A_{i})\setminus B_{i})=m(\sigma_{i}(A_{i})\setminus f(A_{i}))=0$.
Assume for contradiction that there exists an $i$ where $m(\sigma_{i}(A_{i})\setminus f(A_{i}))>0$. Together with $m(f(A_{i})\setminus\sigma_{i}(A_{i}))=0$, it implies that $m(A_{i})=m(\sigma_{i}(A_{i}))>m(f(A_{i}))$. Since we can decompose $\bar{A}$ in terms of disjoint unions of $A_{j}$ where $m(A_{j})\geq m(f(A_{j}))$, it implies that $m(\bar{A})>m(f(\bar{A}))$ and $m(\bar{A}\setminus f(\bar{A}))>0$.
However, let $C=\bar{A}\setminus f(\bar{A})$. Since $\bar{C}\subset\bar{A}$ and $f^{-1}(C)\cap\bar{A}=\emptyset$, we have $f^{-1}(C)\cap\bar{C}=\emptyset$. Let $D=f^{-1}(C)$. Note that $D\cap\bar{C}=\emptyset$ implies that $\bar{D}\cap\bar{C}=\emptyset$. Then, $m(f(D)\setminus\bar{D})=m(C\setminus\bar{D})=m(C)>0$. This is a contradiction. ∎
Proof of Lemma 3.
Because $f$ is not mass-preserving, there exists a measurable set $A\subset\operatorname{\mathbb{R}}^{n}$ such that $m(A)=0$ and $m(f(A))>0$. Thus, we can always find an arbitrarily small closed box $C=\prod_{i=1}^{n}[x_{i},y_{i}]$ where $x_{i},y_{i}]$ and $x_{j},y_{j}]$ are either disjoint or equal such that $m(C\cap f(A))>0$.
Consider the following probability density function
$p(x)=\sum_{i=1}^{n}\frac{\mathbb{1}[x_{i},y_{i}]}{n|y_{i}-x_{i}|}.$
By the definition of locally-dense, we can find an $\alpha\in(0,1]$ and a distribution $G\in\Omega_{0}$ with density function $g$ such that $g(x)|_{\bigcup_{i=1}^{n}[x_{i},y_{i}]}=\alpha p(x)$. Let $X_{1},\cdots,X_{n}$ be random variable with probability density $g$.
Let $X=(X_{1},\cdots,X_{n})$. Then
$\operatorname{\mathbb{P}}[X\in C\cap f(A)]=\frac{m(C\cap f(A))}{m(C)}\operatorname{\mathbb{P}}[X\in C]>0=\operatorname{\mathbb{P}}[X\in f^{-1}(C)\cap A],$
where the last equality is due to the fact that $m(A)=0$. Clearly, $X$ is not invariant under $f$. ∎
Proof of Lemma 4.
Let $n$ be a positive integer. A partition of $n$ is an sequence $\lambda=(\lambda_{1},\lambda_{2},\cdots)$ of non-negative integers in decreasing order
$\lambda_{1}\geq\lambda_{2}\geq\cdots$
such that $\lambda_{1}+\lambda_{2}+\cdots=n$. Let $P_{n}$ denotes the set of all partitions of $n$. We will use $L_{n}$ to denote the reverse lexicographic ordering on the set $P_{n}$ where $L_{n}$ is the subset of $P_{n}\times P_{n}$ consisting of all $(\lambda,\mu)$ such that either $\lambda=\mu$ or the first non-vanishing difference $\lambda_{i}-\mu_{i}$ is positive. $L_{n}$ is a total ordering. We prove the lemma by induction on the reverse lexicographic ordering on partitions of $n$.
Since $f$ is not box-orbit-preserving, there exists a box $A=\prod_{i=1}^{n}[u_{i},v_{i}]$ such that
$m(f(A)\setminus\bar{A})>0.$
Partition $A$ into finitely many disjoint smaller boxes of the form $\prod_{i=1}^{n}[x_{i},y_{i}]$, where each interval $x_{i},y_{i}]$ is either disjoint from or equal to another. Since $m(f(A)\setminus\bar{A})>0$, there exists at least one such box $B$ with
$m(f(B)\setminus\bar{B})>0.$
Each such box $B=\prod_{i=1}^{n}[x_{i},y_{i}]$ induces a partition $\lambda\in P_{n}$ by recording the multiplicities of the intervals $x_{i},y_{i}]$. We proceed by induction on $P_{n}$ under the reverse lexicographic order.
Base case: Suppose $\lambda=(n)$, so $B=[x_{1},y_{1}]^{n}$. Let $C=f(B)\setminus\bar{B}$. Since $m(C)>0$ and $f$ is mass-preserving, $m(f^{-1}(C))>0$.
Consider the density
$p(x)=\frac{1}{|y_{1}-x_{1}|}\mathbf{1}_{[x_{1},y_{1}]}(x).$
By the locally-dense property, there exist $\alpha\in(0,1]$ and a distribution $G\in\Omega_{0}$ with density $g$ such that $g(x)=\alpha p(x)$ on $x_{1},y_{1}]$. Let $X=(X_{1},\dots,X_{n})$ where the $X_{i}$ are drawn i.i.d. from $G$. Then
$\mathbb{P}(X\in f^{-1}(C))=\frac{m(f^{-1}(C))}{m(B)}\alpha^{n}>0,\quad\text{and}\quad\mathbb{P}(X\in C)=0,$
so $X$ is not invariant under $f$.
Inductive step: Assume the statement holds for all partitions $\mu\succ\lambda$.
Let $B=\prod_{i=1}^{n}[x_{i},y_{i}]$ correspond to $\lambda$, and rename the intervals so that $a_{i},b_{i}]$ appears $\lambda_{i}$ times. Define the set
$S=\left\{D=\prod_{i=1}^{n}[w_{i},z_{i}]:[w_{i},z_{i}]\subset\{[a_{1},b_{1}],\dots,[a_{m},b_{m}]\}\right\}\setminus\bar{B},$
and let $S_{\leq}\subset S$ denote those boxes corresponding to partitions less than or equal to $\lambda$ and let $S_{>}\subset S$ denote those boxes corresponding to partitions larger than $\lambda$.
We consider three cases:
Case 1: Suppose there exists $D\in S_{>}$ such that $m(f(B)\cap D)>0$.
Then by the induction hypothesis, there exists a distribution $X$ (with i.i.d. coordinates from a distribution in $\Omega_{0}$) such that $X$ is not invariant under $f$. This is because $\bar{B}\cap\bar{D}=\emptyset$ and $m(f^{-1}(D)\setminus\bar{D})\geq m(f^{-1}(D)\cap B)>0$.
Case 2: Suppose there exists $D\in S_{\leq}$ such that $m(f(B)\cap D)>0$. Let $\mu_{i}$ denotes the number of times $a_{i},b_{i}]$ appears in $D$. Note that $(\mu_{1},\cdots,\mu_{m})$ might not be a partition, but the first non-vanishing difference of $\lambda_{i}-\mu_{i}$ must be positive. Let $j$ be the smallest integer such that $\lambda_{j}>\mu_{j}$.
Let $(\mu_{1},\dots,\mu_{m})$ be the corresponding multiplicities for $D$, and let $j$ be the smallest index such that $\lambda_{j}>\mu_{j}$. Define the density
$p(x)=\sum_{i=1}^{m}(\epsilon_{i-1}-\epsilon_{i})\frac{\mathbf{1}_{[a_{i},b_{i}]}(x)}{|b_{i}-a_{i}|},$
where $\epsilon_{0}=1$, $\epsilon_{m}=0$, and the $\epsilon_{i}$ are chosen later.
By the locally-dense property, there exist $\alpha\in(0,1]$ and $G\in\Omega_{0}$ with density $g$ such that $g(x)=\alpha p(x)$ on $\bigcup_{i}[a_{i},b_{i}]$. Let $X=(X_{1},\dots,X_{n})$ be i.i.d. samples from $G$.
Let $E=f(B)\cap D$. Then $m(f^{-1}(E))>0$, and
$\mathbb{P}(X\in E)\leq\operatorname{\mathbb{P}}(X\in D)=\alpha^{n}\prod_{i=1}^{m}(\epsilon_{i-1}-\epsilon_{i})^{\mu_{i}},$
$\mathbb{P}(X\in f^{-1}(E))=\frac{m(f^{-1}(E))}{m(B)}\alpha^{n}\prod_{i=1}^{m}(\epsilon_{i-1}-\epsilon_{i})^{\lambda_{i}}.$
We choose $\epsilon_{i-1}-\epsilon_{i}=\delta$ small for all $i>j$, and choose $\epsilon_{j-1}-\epsilon_{j}$ so that
$\epsilon_{j-1}-\epsilon_{j}>\frac{m(B)}{m(f^{-1}(E))}\delta.$
Thus,
$\frac{\mathbb{P}(X\in f^{-1}(E))}{\mathbb{P}(X\in E)}\geq\left(\frac{\epsilon_{j-1}-\epsilon_{j}}{\delta}\right)^{\lambda_{j}-\mu_{j}}\frac{m(f^{-1}(E))}{m(B)}>1,$
so $X$ is not invariant under $f$.
Case 3: Suppose $m\left(f(B)\cap S^{c}\cap\bar{B}^{c}\right)>0.$
Define
$p(x)=\sum_{i=1}^{m}\frac{1}{m|a_{i}-b_{i}|}\mathbf{1}_{[a_{i},b_{i}]}(x),$
and apply the locally-dense property to find $\alpha\in(0,1]$ and $G\in\Omega_{0}$ such that $g(x)=\alpha p(x)$ on $\bigcup_{i}[a_{i},b_{i}]$. Let $X=(X_{1},\dots,X_{n})$ be i.i.d. samples from $G$. Let $F=f(B)\cap S^{c}\cap\bar{B}^{c}$.
$\mathbb{P}(X\in F)\leq\operatorname{\mathbb{P}}(X\in S^{c}\cap\bar{B}^{c})=0,$
but
$\mathbb{P}(X\in f^{-1}(F))=\frac{m(f^{-1}(F))}{m(B)}\mathbb{P}(X\in B)>0,$
so again $X$ is not invariant under $f$. This concludes the proof. ∎
B.3 Statements and proofs of claims in Remark 4
Proposition 9.
The statement in Proposition 7 holds if we replace $\Omega_{\text{cts}}(\mathbb{R})$ with $\Omega_{\text{cts}}([b_{0},b_{1}])$ for some interval $b_{0},b_{1}]$ and $\beta$ in the support. The same holds with Proposition 8 for any $q\in(b_{0},b_{1})$.
Proof of Proposition 9.
We first consider the statement in Proposition 7 replacing $\Omega_{\text{cts}}(\mathbb{R})$ with $\Omega_{\text{cts}}([b_{0},b_{1}])$ for some interval $b_{0},b_{1}]$. We will separately consider the cases where $\beta$ is odd and even.
For the case where $\beta$ is odd, consider any positive integer $k$, and consider $\beta$ such that $b_{0}^{2k+1}<\beta<b_{1}^{2k+1}$. For any $m$, we will show that the null is locally-dense by considering intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset\mathcal{B}$ where
$\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\min\{b_{1}-\beta^{\frac{1}{2k+1}},\beta^{\frac{1}{2k+1}}-b_{0}\}}{m+1}.$
Consider any $X$ with probability density $p$ on such the intervals. We want to show that we can find an $\alpha\in(0,1]$ and a probability density function $g$ on $\mathcal{B}$ with $2k+1$-th moment equal to $\beta$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$. To do so, let $h$ be a probability density function supported on $\mathcal{B}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=ap(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then
$\operatorname{\mathbb{E}}[X_{a}^{2k+1}]=a\operatorname{\mathbb{E}}[X^{2k+1}]+(1-a)\operatorname{\mathbb{E}}[Y^{2k+1}].$
If $\operatorname{\mathbb{E}}[X^{2k+1}]\geq\beta$, given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\beta^{\frac{1}{2k+1}}-b_{0}}{m+1}$, we can find an interval $x,y]\subset[b_{0},\beta^{\frac{1}{2k+1}}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Let
$h(x)=\frac{1}{y-x}\mathbb{1}_{[x,y]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k+1}]\leq y^{2k+1}\leq\beta.$
We can then find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k+1}_{\alpha}]=\beta$.
Similarly, if $\operatorname{\mathbb{E}}[X^{2k+1}]\leq\beta$, given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{b_{1}-\beta^{\frac{1}{2k+1}}}{m+1}$, we can find an interval $x,y]\subset[\beta^{\frac{1}{2k+1}},b_{1}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Let
$h(x)=\frac{1}{y-x}\mathbb{1}_{[x,y]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k+1}]\geq x^{2k+1}\geq\beta.$
We can then find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k+1}_{\alpha}]=\beta$.
For the case where $\beta$ is even, we will show that for any real number $0<\beta<\max\{b_{0}^{2k},b_{1}^{2k}\}$ and any continuous random variable $X$ with probability density function $p$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset\mathcal{B}$ where
$\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\min\{\max\{|b_{0}|,|b_{1}|\}-\beta^{\frac{1}{2k}},\beta^{\frac{1}{2k}}\}}{m+1},$
we can find an $\alpha\in(0,1]$ and a probability density function $g$ on $\mathcal{B}$ with $2k$-th moment equal to $\beta$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$.
To show this, let us assume without loss of generality that $|b_{1}|\geq|b_{0}|$. Let $h$ be a probability density function supported on $\mathcal{B}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=ap(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then
$\operatorname{\mathbb{E}}[X_{a}^{2k}]=a\operatorname{\mathbb{E}}[X^{2k}]+(1-a)\operatorname{\mathbb{E}}[Y^{2k}].$
If $\operatorname{\mathbb{E}}[X^{2k}]\geq\beta$, given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\beta^{\frac{1}{2k}}}{m+1}$, we can find an interval $x,y]\subset[-\beta^{\frac{1}{2k}},\beta^{\frac{1}{2k}}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Let
$h(x)=\frac{1}{y-x}\mathbb{1}_{[x,y]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k}]\leq y^{2k}<\beta.$
We can then find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k}_{\alpha}]=\beta$.
Similarly, if $\operatorname{\mathbb{E}}[X^{2k}]\leq\beta$, given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{|b_{1}|-\beta^{\frac{1}{2k}}}{m+1}$, we can find an interval $x,y]\subset[\beta^{\frac{1}{2k}},\mathbb{1}\left[b_{1}>0\right]\times b_{1}-\mathbb{1}\left[b_{1}<0\right]\times b_{1}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Let
$h(x)=\frac{1}{y-x}\mathbb{1}_{[x,y]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k}]\geq x^{2k}\geq\beta.$
We can then find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k}_{\alpha}]=\beta$.
We next consider the statement in Proposition 8 replacing $\Omega_{\text{cts}}(\mathbb{R})$ with $\Omega_{\text{cts}}([b_{0},b_{1}])$ for some interval $b_{0},b_{1}]$ and considering any $q\in(b_{0},b_{1})$. Let $X$ be a continuous random variable with probability density function $f$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset[b_{0},b_{1}]$ where $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\min\{\frac{q-b_{0}}{m+1},\frac{b_{1}-q}{m+1}\}$. We will show that for any $q\in(0,1)$, we can find an $\alpha\in(0,1]$ and a probability density function $g$ on the real line with the $p$-th quantile equal to $q$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha f(x)$.
Let $h$ be a probability density function supported on $b_{0},b_{1}]\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=af(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then,
$\operatorname{\mathbb{P}}[X_{a}\leq q]=a\operatorname{\mathbb{P}}[X\leq q]+(1-a)\operatorname{\mathbb{P}}[Y\leq q].$
Given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\min\{\frac{q-b_{0}}{m+1},\frac{b_{1}-q}{m+1}\}$, we can find intervals $x,y]\subset[b_{0},q]$ and $w,z]\subset(q,b_{1}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=[w,z]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Now we can choose $\alpha=a>0$ small enough such that $\alpha\operatorname{\mathbb{P}}[X\leq q]<p$. In order for $\operatorname{\mathbb{P}}[X_{\alpha}\leq q]=p$, we need $\operatorname{\mathbb{P}}[Y\leq q]=\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}.$ This can be achieved by letting $h(x)=\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{(1-\alpha)(y-x)}\mathbb{1}_{[x,y]}+(1-\frac{p-a\operatorname{\mathbb{P}}[X\leq q]}{(1-\alpha)})\frac{1}{(z-w)}\mathbb{1}_{[w,z]}$. ∎
Proposition 10.
For any $\beta\in\mathbb{R}_{>0}$, $\Omega_{0}=\{P\in\Omega_{\text{cts}}(\mathbb{R}):\operatorname{Var}_{P}[X_{i}]=\beta\}$ is locally-dense and thus does not satisfy the randomization hypothesis.
Proof of Proposition 10.
Let $X$ be a continuous random variable with probability density function $p$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]$. For any $\beta\in\operatorname{\mathbb{R}}_{>0}$, we will show that we can find an $\alpha\in(0,1]$ and a probability density function $g$ on the real line with variance $\beta$ such that $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$.
Let $h$ be a probability density function supported on $\operatorname{\mathbb{R}}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=ap(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$.
Then
$\begin{split}\operatorname{Var}[X_{a}]&=a(\operatorname{Var}[X]+\operatorname{\mathbb{E}}[X]^{2})+(1-a)(\operatorname{Var}[Y]+\operatorname{\mathbb{E}}[Y]^{2})-(a\operatorname{\mathbb{E}}[X]+(1-a)\operatorname{\mathbb{E}}[Y])^{2}\\ &=a[(1-a)\operatorname{\mathbb{E}}[X]^{2}+\operatorname{Var}[X]+(1-a)\operatorname{\mathbb{E}}[Y]^{2}-2(1-a)\operatorname{\mathbb{E}}[X]\operatorname{\mathbb{E}}[Y]]+(1-a)\operatorname{Var}[Y].\end{split}$
Choose $M>0$ such that $\bigcup_{i=1}^{m}[x_{i},y_{i}]\subset[-M,M]$. For any $\beta>0$ and $a=\frac{1}{2}$, we can find $z>M+1$ such that
$a[(1-a)\operatorname{\mathbb{E}}[X]^{2}+\operatorname{Var}[X]+(1-a)z^{2}-2(1-a)z\operatorname{\mathbb{E}}[X]]>\beta.$
We then choose $w\in(0,1)$ such that $\frac{1}{12}w^{2}<\beta$. Let $h(x)=\frac{1}{w}\mathbb{1}_{[z-\frac{1}{2}w,z+\frac{1}{2}w]}$. In this case, since $\operatorname{Var}[Y]\geq 0$, we have $\operatorname{Var}[X_{1/2}]>\beta$. On the other hand, $\operatorname{Var}[X_{0}]=\operatorname{Var}[Y]=\frac{1}{12}w^{2}<\beta$. Because $\operatorname{Var}[X_{a}]$ is a continuous in $a$, we can find an $\alpha\in(0,\frac{1}{2})$ such that $\operatorname{Var}[X_{\alpha}]=\beta$. Thus, for any $\beta\in\operatorname{\mathbb{R}}_{>0}$, $\{P\in\Omega_{\text{cts}}:\operatorname{Var}_{P}[X_{i}]=\beta\}$ is locally-dense. ∎