### v2

Abstract
Randomization tests deliver exact finite-sample Type 1 error control when the null satisfies the randomization hypothesis. In practice, achieving these guarantees often requires stronger conditions than the null hypothesis of primary interest. For example, sign-change tests of mean zero require symmetry and need not control finite-sample size for non-symmetric mean-zero distributions. We investigate whether the mismatch between the null and the invariance conditions required for exactness reflects the use of particular transformations or a more fundamental limitation. We provide a simple necessary and sufficient condition for a null hypothesis to admit a randomization test. Applying this framework to one-sample problems, we characterize the nulls that admit randomization tests on finite supports and derive impossibility results on continuous supports. In particular, we show that several common nulls, including mean zero, do not admit randomization tests. We further show that, among one-sample tests using linear group actions, the admissible nulls are limited to subsets of symmetric or Gaussian distributions. These results confirm that the absence of exact finite-sample validity is inherent for many commonly studied nulls and that practitioners using existing tests are not foregoing feasible exact alternatives.
1 Introduction
Consider the problem of testing a null hypothesis given a finite sample of data. If this null satisfies a group invariance property referred to as the “randomization hypothesis,” then one can construct randomization tests that obtain exact finite-sample Type 1 error control. Intuitively, this property requires that, under the null, certain transformations of the data leave its distribution unchanged. In part due to the appeal of finite-sample validity, there is a large methodological literature on randomization tests and a similarly large applied literature using these methods (for a review, see Ritzwoller et al. 2024).
In practice, randomization tests often rely on invariance conditions that are stronger than the null hypothesis of primary interest. This is a problem because it re-introduces an inability to control Type 1 error: the test can over-reject when the data is drawn from a distribution that satisfies the hypothesis of primary interest but not the stronger conditions. For example, a commonly used randomization test for the null of mean zero with i.i.d. observations is based on sign changes. This test controls finite-sample size only under symmetry and may over-reject for non-symmetric mean-zero distributions.
This paper investigates whether the failure to control Type 1 error for certain null hypotheses (e.g. the null of mean zero) is due to the specific randomization tests used in practice (e.g. the use of sign changes) or from a more fundamental limitation. To formalize this distinction, we develop a framework that characterizes which null hypotheses admit randomization tests. We provide a simple necessary and sufficient condition for a null to satisfy the randomization hypothesis that avoids group-theoretic structure.
We then apply this framework to one-sample problems. We show that certain null hypotheses—such as the null of mean zero—do not admit randomization tests that achieve exact finite-sample validity. These results relate to classical non-existence results for nonparametric inference (Bahadur and Savage 1956, Lehmann and Loh 1990, Romano 2004). By focusing on randomization tests, we derive conditions for the existence and non-existence of exact finite-sample procedures across a range of null hypotheses and distributional classes. These results imply that existing applications of randomization tests, such as sign-change tests, are not overlooking alternative procedures that would achieve exact finite-sample validity.
We also provide guidance for constructing randomization tests when standard approaches are not applicable. On finite supports, we give an explicit characterization of nulls that admit randomization tests. On continuous supports, we study tests based on linear group actions. We show that, in one-sample settings, such tests are very limited: the admissible nulls correspond to subsets of symmetric or Gaussian distributions.
As part of identifying whether a null admits a randomization test, we also show how our framework can be used to construct these tests. We do so by providing explicit examples — including a construction of a normality test that makes use of rotation symmetries — that may be of theoretic or applied interest.
Taken together, these results yield a positive implication for applied work: the use of standard randomization tests does not overlook alternative procedures that achieve exact finite-sample validity for many commonly studied nulls. They also affirm and motivate a focus on asymptotic properties of randomization tests (see, e.g., Romano (1989, 1990), DiCiccio and Romano (2017), Canay et al. (2017), Lei and Bickel (2021), Pouliot (2024), Bai et al. (2024)). Finally, our results suggest that one-sample randomization tests that go beyond symmetry or normality must rely on non-linear transformations.
The remainder of the paper is organized as follows. Section 2 reviews the randomization test procedure and develops our framework. Section 3 studies one-sample tests with finite support. Section 4 considers continuous supports. Section 5 concludes.
2 The randomization test
2.1 Construction of the randomization test
The construction follows Lehmann et al. (1986); see Zhang and Zhao (2023) for a recent formulation. Let $X=(X_{1},\dots,X_{n})$ be data taking values in $\mathcal{X}$ with distribution $P\in\Omega$. Consider the problem of testing $H_{0}:P\in\Omega_{0}$, and suppose that there is a group $\mathbb{G}$ of measurable transformations $g:\mathcal{X}\to\mathcal{X}$ such that $gX\stackrel{{\scriptstyle d}}{{=}}X$ whenever $X\sim P\in\Omega_{0}$. For ease of exposition, we assume $\mathbb{G}$ is finite with size $M$.
Let $T:\mathcal{X}\to\mathbb{R}$ be any test statistic used to test the null. For every $x\in\mathcal{X}$, let $T^{(1)}(x;\mathbb{G})\leq\dots\leq T^{(M)}(x;\mathbb{G})$ be the ordered values of $T(gx)$ as $g$ ranges over $\mathbb{G}$. Given nominal level $\alpha\in(0,1)$, let $k\equiv M-\lfloor M\alpha\rfloor,$ where $\lfloor M\alpha\rfloor$ denotes the largest integer less than or equal to $M\alpha$. Let $M^{+}(x)$ and $M^{0}(x)$ be the number of values $T^{(j)}(x)$ that are greater than $T^{(k)}(x)$ and equal to $T^{(k)}(x)$, respectively. Set
$a(x)=\frac{M\alpha-M^{+}(x)}{M^{0}(x)}.$
Define the randomization test as
$\phi(x;\mathbb{G})\equiv\begin{cases}1&\text{if $T(x)>T^{(k)}(x;\mathbb{G})$}\\ a(x)&\text{if $T(x)=T^{(k)}(x;\mathbb{G})$}\\ 0&\text{if $T(x)<T^{(k)}(x;\mathbb{G})$}\end{cases},$ (1)
The following Theorem shows that the randomization test has nominal level $\alpha$. The proof we present is slightly different than the proof in Lehmann et al. (1986), in part due to our focus on the role of the group structure $\mathbb{G}$.
Theorem 1.
Suppose that for any $g\in\mathbb{G}$, $gX\stackrel{{\scriptstyle d}}{{=}}X$ whenever $X\sim P\in\Omega_{0}$. Then $\operatorname{\mathbb{E}}_{P}[\phi(X;\mathbb{G})]=\alpha$ $\forall\;P\in\Omega_{0}$.
Proof of Theorem 1.
Consider any $X\sim P\in\Omega_{0}.$ Then
$\operatorname{\mathbb{E}}_{P}[\phi(X;\mathbb{G})]=\frac{1}{M}\operatorname{\mathbb{E}}_{P}\left[\sum_{g\in\mathbb{G}}\phi(X;\mathbb{G})\right]=\frac{1}{M}\operatorname{\mathbb{E}}_{P}\left[\sum_{g\in\mathbb{G}}\phi(gX;g\mathbb{G})\right],$
where the last equality used the form of $\phi$ as in (1) and that $X\stackrel{{\scriptstyle d}}{{=}}gX$ for all $g\in\mathbb{G}$. By construction, $\frac{1}{M}\sum_{g\in\mathbb{G}}\phi(gx;\mathbb{G})=\alpha$ for any $x\in\mathcal{X}$, so that $\frac{1}{M}\operatorname{\mathbb{E}}_{P}\left[\sum_{g\in\mathbb{G}}\phi(gX;\mathbb{G})\right]=\alpha.$ To complete the proof, it thus suffices to show that
$\operatorname{\mathbb{E}}_{P}\left[\sum_{g\in\mathbb{G}}\phi(gX;\mathbb{G})\right]=\operatorname{\mathbb{E}}_{P}\left[\sum_{g\in\mathbb{G}}\phi(gX;g\mathbb{G})\right].$ (2)
This follows from the group property that $g\mathbb{G}=\mathbb{G}$ for all $g\in\mathbb{G}$, and thus $\phi(gx;\mathbb{G})=\phi(gx;g\mathbb{G})$ for all $x$ and $g\in\mathbb{G}$, so that (2) holds. ∎
Remark 1.
Suppose that we did not require that the collection of invertible transformations $\mathbb{G}$ is a group. For (2) to hold generally, we require that $\mathbb{G}g=\mathbb{G}$ for all $g\in\mathbb{G}$. A basic property of groups is that if the set $\mathbb{G}$ is a collection of invertible transformations, then $\mathbb{G}g=\mathbb{G}$ for all $g\in\mathbb{G}$ if and only if $\mathbb{G}$ is a group (with composition as the operation). In this sense, the requirement that $\mathbb{G}$ is a group is both necessary and sufficient.
2.2 Framework
Having established the construction, we now develop a simple framework to examine which null hypotheses admit a group that is invariant to all distributions in the null.
Definition 1 (Randomization hypothesis).
We say $\Omega_{0}$ satisfies the randomization hypothesis if there exists a group $\mathbb{G}$ of measurable transformations $g:\mathcal{X}\to\mathcal{X}$ such that
$\Omega_{0}\subseteq\Omega_{\mathbb{G}}\subset\Omega,$ (3)
where $\Omega_{\mathbb{G}}\equiv\{P\in\Omega\;:\;gX\stackrel{{\scriptstyle d}}{{=}}X\;\forall\;g\in\mathbb{G}\ ,\ X\sim P\}.$
Remark 2.
The set inclusions in (3) enforce that $\Omega_{0}$ is invariant to $\mathbb{G}$ while also ensuring that $\mathbb{G}$ is not invariant to all distributions in $\Omega$. The latter condition rules out trivial groups (e.g. the identity) that are invariant to all distributions.
Our analysis will focus on characterizing which null hypotheses do and do not satisfy the randomization hypothesis. For those that do, we will be interested in constructing a group to implement the test. The following result provides a tool to answer both of these questions in a way that avoids group-theoretic concepts.
Proposition 1.
$\Omega_{0}$ satisfies the randomization hypothesis if and only if there exists a bijective measurable function $f:\mathcal{X}\to\mathcal{X}$ such that
$\Omega_{0}\subseteq\Omega_{f}\subset\Omega,$ (4)
where $\Omega_{f}\equiv\big\{P\in\Omega\;:\;fX\stackrel{{\scriptstyle d}}{{=}}X,X\sim P\big\}.$ When this holds, a randomization test can be constructed by using the group generated by $f$.
All omitted proofs can be found in Supplementary Appendix B. The proof of the forward direction is immediate and the proof of the reverse direction follows from showing that if $f$ is invariant to all distributions in $\Omega_{0}$, then so is $f^{-1}$. We can then show that $\Omega_{0}$ satisfies the randomization hypothesis using the group generated by $f$.
Proposition 1 shows that to study which null hypotheses satisfy the randomization hypothesis, it suffices to consider individual bijective functions. Given such a function satisfying (4), a randomization test can be implemented using the group it generates.
In the remainder of the paper, we study one-sample problems with $X=(X_{1},\dots,X_{n})$ with $X_{i}$ i.i.d. from $P\in\Omega$, supported on a subset of the real line. Here, $\Omega$ denotes the set of distributions for $X_{i}$.
3 One-sample tests on a finite support
In this section, we suppose $X_{i}$ are drawn i.i.d. from $P$ supported on a finite set $\mathcal{A}\equiv\{\alpha_{1},\dots,\alpha_{K}\}\subset\mathbb{R}$. We represent $P$ by its probability mass function $p:\mathcal{A}\to[0,1]$ and define
$\Omega_{\text{discrete}}\equiv\left\{p:\mathcal{A}\to[0,1]\text{ s.t. }\sum_{k=1}^{K}p(\alpha_{k})=1\right\}.$ (5)
By Proposition 1, a null hypothesis $\Omega_{0}\subset\Omega_{\text{discrete}}$ satisfies the randomization hypothesis if and only if there exists a bijection $f:\mathcal{A}^{n}\to\mathcal{A}^{n}$ such that $\Omega_{0}\subseteq\Omega_{f}\subset\Omega_{\text{discrete}}$. On a finite support, we can explicitly write $\Omega_{f}$ as
$\Omega_{f}=\bigcap_{x\in\mathcal{A}^{n}}\left\{p\in\Omega_{\text{discrete}}:\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(f_{j}(x))\right\}.$ (6)
Each of the individual sets in this intersection corresponds to the invariances of simpler functions that equal $f$ for a single value and otherwise equal the identity. This provides an explicit characterization of Proposition 1 for one-sample tests over finite supports.
Proposition 2.
$\Omega_{0}\subset\Omega_{\text{discrete}}$ satisfies the randomization hypothesis if and only if there exists $x,y\in\mathcal{A}^{n}$ such that $y$ is not a permutation of $x$ and such that $\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(y_{j})$ for all $p\in\Omega_{0}$.
Proposition 2 is useful for two reasons. First, it can be used to identify bijective functions that satisfy the invariance. We can then construct randomization tests by taking the groups generated by the functions. To illustrate this, Example 1 of Supplementary Appendix A uses the result to construct the usual test of symmetry around zero and to construct a test of two points having equal mass.
Second, Proposition 2 shows that null hypotheses satisfying the randomization hypothesis are characterized by equal probability mass assigned to distinct configurations (accounting for multiplicities). Null hypotheses not defined by such restrictions therefore do not satisfy the randomization hypothesis. In what follows, we apply Proposition 2 to show that null hypotheses defined by $k$-th moments (including the mean) and quantiles do not satisfy the randomization hypothesis. These results illustrate how the proposition can be used to determine whether a null hypothesis admits a randomization test.
Proposition 3.
If $|\mathcal{A}|>4$, then for any positive integer $t$ and $\beta\in(\alpha^{t}_{1},\alpha^{t}_{K})$, $\Omega_{0}=\{P\in\Omega_{\text{discrete}}:\operatorname{\mathbb{E}}_{P}[X_{i}^{t}]=\beta\}$ does not satisfy the randomization hypothesis.
Proposition 4.
If $|\mathcal{A}|>2$, then for any $p\in(0,1)$ and $q\in(\alpha_{1},\alpha_{K})$, $\Omega_{0}=\{P\in\Omega_{\text{discrete}}:\operatorname{\mathbb{P}}[X_{i}\leq q]=p\}$ does not satisfy the randomization hypothesis.
The proofs of Propositions 3 and 4 proceed by showing that for any $x,y\in\mathcal{A}^{n}$ that are not equal up to permutation, one can construct a distribution in the null that assigns different probabilities to $x$ and $y$. The result then follows from Proposition 2.
Remark 3.
The cardinality requirements in Propositions 3 and 4 exclude edge cases. For example, when $|\mathcal{A}|=3$ and $\mathcal{A}=\{-L,0,L\}$, the null of mean zero coincides with symmetry, which satisfies the randomization hypothesis via sign changes. The proof of Proposition 3 provides an analogous edge case for $|\mathcal{A}|=4$. For Proposition 4, the result continues to hold when $|\mathcal{A}|=2$ except at $p=1/2$, where sign changes yield a valid randomization test.
4 One-sample tests on a continuous support
In this section, we suppose $X_{i}$ is drawn i.i.d. from $P$ supported on the real line. We first characterize all null hypotheses that are invariant to linear group actions. We then consider arbitrary groups and develop a necessary condition for a null hypothesis to satisfy the randomization hypothesis.
4.1 Groups of linear transformations
A group of linear transformations $\mathbb{G}$ is one in which each element $g$ corresponds to a linear map $A_{g}:\mathbb{R}^{n}\to\mathbb{R}^{n}$. The following result characterizes the distributions invariant to such groups.
Proposition 5.
Let $P\in\Omega_{\text{cts}}$, where $\Omega_{\text{cts}}$ is a subset of distributions supported on the real line. Let $\mathbb{G}$ be a group of linear transformations acting on $\mathbb{R}^{n}$. Then exactly one of the following is true: 1. $\Omega_{\mathbb{G}}=\Omega_{\text{cts}}$, 2. $\Omega_{\mathbb{G}}=\{P\in\Omega_{\text{cts}}:\text{$P$ symmetric about $0$}\}$, 3. $\Omega_{\mathbb{G}}\subseteq\{N(\mu,\sigma^{2})\in\Omega_{\text{cts}}:\mu\in\mathbb{R},\sigma^{2}>0\}$, or 4. $\Omega_{\mathbb{G}}=\emptyset$.
The proof proceeds by writing $gX=AX$ for some matrix $A$. If $A$ is diagonal, invariance requires each coordinate to be symmetric about zero. If $A$ is not diagonal, then at least one component of $AX$ is a non-trivial linear combination of multiple coordinates of $X$. Since both $AX$ and $X$ consist of i.i.d. components, the Darmois-Skitovich theorem implies that $X_{i}$ must be Gaussian.
Proposition 5 shows that, when restricted to linear group actions, one-sample randomization tests extend permutation-based invariance in limited ways. In particular, they can test membership in the class of distributions symmetric about zero or in subsets of the Gaussian family. The former can be implemented using sign changes. Example 2 in Supplementary Appendix A illustrates the latter.
4.2 Arbitrary groups
We now allow $\mathbb{G}$ to be arbitrary. The classical result of Bahadur and Savage (1956) shows that sufficiently rich classes of distributions cannot be tested in a nonparametric setting. In our context, a related conclusion holds under a local version of this richness condition.
Let $\Omega_{\text{cts}}(\mathcal{B})$ be the set of distributions supported on $\mathcal{B}\subseteq\mathbb{R}$. Informally, a null is locally-dense if it can locally replicate any density on finitely many small intervals (up to scale).
Definition 2 (Locally-dense).
A set $\Omega_{0}\subset\Omega_{\text{cts}}(\mathcal{B})$ is locally-dense if there exists $L:\mathbb{N}\to\mathbb{R}_{+}$ such that for any collection of intervals $I_{1},\dots,I_{m}$ with $|I_{i}|<L(m)$ and any density $p$ supported on $\bigcup_{i=1}^{m}I_{i}$, there exist $H\in\Omega_{0}$ with density $h$ and $\alpha\in(0,1]$ such that
$h(x)=\alpha p(x)\quad\text{for all }x\in\bigcup_{i=1}^{m}I_{i}.$
The following result shows that locally-dense null hypotheses do not satisfy the randomization hypothesis.
Proposition 6.
If $\Omega_{0}$ is locally-dense, then $\Omega_{0}$ does not satisfy the randomization hypothesis.
The proof uses Proposition 1 and shows that for any function $f$ with $\Omega_{0}\subseteq\Omega_{f}$, local denseness implies $\Omega_{f}=\Omega$. Intuitively, local denseness ensures that the null can approximate any density on finitely many regions, allowing the invariance to extend to all distributions.
We now use Proposition 6 to show that the same negative results obtained in Section 3 also hold when considering continuous distributions or when considering continuous distributions over a fixed interval.
Proposition 7.
For any positive integer $t$ and $\beta\in\mathbb{R}$, $\Omega_{0}=\{P\in\Omega_{\text{cts}}(\mathbb{R}):\operatorname{\mathbb{E}}_{P}[X_{i}^{t}]=\beta\}$ is locally-dense and thus does not satisfy the randomization hypothesis.
To illustrate the content of this result, consider the null $\operatorname{\mathbb{E}}_{P}[X_{i}]=\beta$. Let $I_{1},\dots,I_{m}$ be intervals with width bounded by $\delta>0$. For any density $p$ supported on these intervals with mean $\gamma_{p}$ and any $\alpha\in(0,1]$, construct a distribution $H$ with density $\alpha p(x)$ on these intervals. Then $H$ has mean
$\alpha\gamma_{p}+(1-\alpha)\operatorname{\mathbb{E}}_{H}[X_{i}\mid X_{i}\notin\cup_{j}I_{j}].$
By placing the remaining mass on two intervals above and below $\beta$, one can choose $\alpha$ and the remaining mass so that the mean equals $\beta$. Thus the null is locally-dense. The following results all make use of similar constructions.
Proposition 8.
For any $p\in(0,1)$ and $q\in\operatorname{\mathbb{R}}$, $\Omega_{0}=\{P\in\Omega_{\text{cts}}(\mathbb{R}):\operatorname{\mathbb{P}}[X_{i}\leq q]=p\}$ is locally-dense and thus does not satisfy the randomization hypothesis.
Remark 4.
We show in Supplementary Appendix B.3 that Proposition 7 extends to $\Omega_{\text{cts}}([b_{0},b_{1}])$ for interval $b_{0},b_{1}]$ and $\beta$ in the support. The same extension holds for Proposition 8 for any $q\in(b_{0},b_{1})$. We also show that the null hypothesis of fixed variance is locally-dense and therefore does not satisfy the randomization hypothesis.
5 Conclusion
Randomization tests yield exact finite-sample Type 1 error control when the null satisfies the randomization hypothesis. In practice, randomization tests often rely on invariance conditions that are stronger than the null hypothesis of primary interest. This paper developed a framework for characterizing which null hypotheses admit randomization tests with exact finite-sample validity.
Applying our framework to one-sample problems revealed impossibilities and limited possibilities of uses of finite-sample randomization tests. Many standard null hypotheses, including the null of mean zero, do not admit randomization tests with exact finite-sample validity. We further show that, when restricted to linear group actions, one-sample randomization tests can only test membership in the class of symmetric distributions or in subsets of the Gaussian family. These results imply that existing procedures are not overlooking alternative randomization tests that would achieve finite-sample validity.
While testing beyond symmetry or normality requires non-linear transformations, we show that even such transformations cannot be used to test many commonly studied null hypotheses. More broadly, the framework developed here delineates the limits of finite-sample randomization tests and provides constructive guidance for identifying when nulls admit such tests and how to construct them when they exist. Overall, our findings indicate that current practice is not omitting feasible exact procedures and further motivate a focus on asymptotic validity.
SUPPLEMENTARY MATERIAL
Appendix A Examples of randomization tests
Example 1 (Construction of randomization tests over discrete support).
  1. 1.
    For $L>0$, let $\mathcal{A}\equiv\{-L,\dots,0,\dots,L\}$ and let $\Omega_{0}$ be the set of distributions that are symmetric around 0. Consider any $x,y$ with $x_{j}\neq 0$ for some $j$ and y=xy=-x. We then have that $c(x_{j};x)>c(x_{j};y)$ and that for all $p\in\Omega_{0}$, $\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(-x_{j})=\prod_{j=1}^{n}p(y_{j})$. By Proposition 2 $\Omega_{0}$ satisfies the randomization test. We can then construct a randomization test for $\Omega_{0}$ by taking the group arising from any collection of functions that send some values of $x\in\mathcal{A}^{n}$ to x-x and otherwise act as the identity function.
  2. 2.
    For any $\mathcal{A}$, let $\Omega_{0}$ be the set of distributions where $p(\alpha_{\ell})=p(\alpha_{\ell^{\prime}})$ for some $\alpha_{\ell}\neq\alpha_{\ell^{\prime}}$. Consider any $x,y$ with $x$ containing $\alpha_{\ell}$ but not $\alpha_{\ell^{\prime}}$ and $y_{j}=\alpha_{\ell^{\prime}}\mathbb{1}[x_{j}=\alpha_{\ell}]+\alpha_{\ell}\mathbb{1}[x_{j}=\alpha_{\ell^{\prime}}]+x_{j}\mathbb{1}[x_{j}\neq\alpha_{\ell},x_{j}\neq\alpha_{\ell^{\prime}}]$. We then have that $c(\alpha_{\ell};x)>c(\alpha_{\ell};y)$ and for all $p\in\Omega_{0}$, $\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(y_{j})$. By Proposition 2 $\Omega_{0}$ satisfies the randomization test. We can then construct a randomization test for $\Omega_{0}$ by taking the group arising from any collection of functions that acts as the identity except for swapping the values of $\alpha_{\ell}$ and $\alpha_{\ell^{\prime}}$.
Example 2 (Randomization tests for Gaussian distributions).
  1. 1.
    Let $\mathbb{G}$ be the orthogonal group of degree $n$. By Maxwell’s Theorem, the set of invariances given $\mathbb{G}$ is $\Omega_{\mathbb{G}}=\{N(0,\sigma^{2}):\mu\in\mathbb{R},\sigma^{2}>0\}.$ Thus, this group can be used to construct a randomization test of a distribution being Gaussian with mean zero.
  2. 2.
    Let $\mathbb{G}$ be the set of transformations generated by $3\times 3$ block permutations of the matrix
    $\begin{pmatrix}\frac{2}{3}&-\frac{1}{3}&\frac{2}{3}\\ \frac{2}{3}&\frac{2}{3}&-\frac{1}{3}\\ -\frac{1}{3}&\frac{2}{3}&\frac{2}{3}\\ \end{pmatrix}$
    In the below proof, we show that $\mathbb{G}$ is a group and that $\Omega_{\mathbb{G}}=\{N(\mu,\sigma^{2}):\mu\in\mathbb{R},\sigma^{2}>0\}.$ Thus, this group can be used to construct a randomization test of a distribution being Gaussian.
Proof of claims in Example 2.
Let $X_{1},X_{2},X_{3}$ be i.i.d. draws from $N(\mu,\sigma^{2})$. Let $X=(X_{1},X_{2},X_{3})$ and $V=(\mu,\mu,\mu)$. Because
$A\equiv\begin{pmatrix}\frac{2}{3}&-\frac{1}{3}&\frac{2}{3}\\ \frac{2}{3}&\frac{2}{3}&-\frac{1}{3}\\ -\frac{1}{3}&\frac{2}{3}&\frac{2}{3}\\ \end{pmatrix}$
is orthogonal, we know from Maxwell’s Theorem,
$A(X-V)\stackrel{{\scriptstyle d}}{{=}}X-V.$
Since $AV=V$, we have that
$AX\stackrel{{\scriptstyle d}}{{=}}X.$
Inductively, $A^{n}X\stackrel{{\scriptstyle d}}{{=}}X$ for all $n\in\mathbb{Z}$. Thus, $\{N(\mu,\sigma^{2}):\mu\in\mathbb{R},\sigma^{2}>0\}\subset\Omega_{\mathbb{G}}$. By proposition 5, $\{N(\mu,\sigma^{2}):\mu,\sigma\in\operatorname{\mathbb{R}}\}=\Omega_{\mathbb{G}}$. ∎
Appendix B Proofs
Appendix B.1 provides the proofs of all propositions except for the proof of Proposition 6. The proof of Proposition 6 requires additional notation and definitions and is developed in Appendix B.2. Lastly, Appendix B.3 provides statements and proofs of the claims made in Remark 4.
B.1 All proofs excluding the proof of Proposition 6
Proof of Proposition 1.
The forward direction follows from noting that $\Omega_{\mathbb{G}}=\cap_{g\in\mathbb{G}}\Omega_{g}$, and since (3) holds, there must be at least some $f\in\mathbb{G}$ such that (4) holds. Group elements have inverses and thus $f$ is bijective.
For the reverse direction, consider a bijective measurable $f$ satisfying (4). Note that the set generated by $f$, denoted $\langle f\rangle\equiv\{f^{K}:K\in\mathbb{Z}\}$, is a group with composition as the operation, with $f^{-K}={f^{-1}}^{K}$ where $f^{-1}$ denotes the inverse of $f$. Then $\Omega_{\langle f\rangle}\subseteq\Omega_{f}\subset\Omega$.
It remains to show that $\Omega_{0}\subseteq\Omega_{\langle f\rangle}$, which is equivalent to showing that $\operatorname{\mathbb{P}}[f^{K}X\in A]=\operatorname{\mathbb{P}}[X\in A]$ for all $X\sim P\in\Omega_{0}$, $K\in\mathbb{Z}$, and measurable $A$. When $K=0$, the proof follows trivially since the identity is invariant to all distributions. For any integer $K>0$,
$\operatorname{\mathbb{P}}[f^{K}X\in A]=\operatorname{\mathbb{P}}[fX\in f^{K-1}A]=\operatorname{\mathbb{P}}[X\in f^{K-1}A]=\operatorname{\mathbb{P}}[f^{K-1}X\in A],$ (7)
where we used that $f$ is invertible and also used that $f$ is measurable so that $f^{K-1}A$ is measurable and thus, since $\Omega_{0}\subseteq\Omega_{f}$, we have $\operatorname{\mathbb{P}}[fX\in f^{K-1}A]=\operatorname{\mathbb{P}}[X\in f^{K-1}A]$. Repeating this step, we obtain $\operatorname{\mathbb{P}}[f^{K}X\in A]=\operatorname{\mathbb{P}}[X\in A]$. Lastly, note that
$\operatorname{\mathbb{P}}[f^{-1}X\in A]=\operatorname{\mathbb{P}}[X\in fA]=\operatorname{\mathbb{P}}[fX\in fA]=\operatorname{\mathbb{P}}[X\in A],$ (8)
where the second equality used that $\Omega_{0}\subseteq\Omega_{f}$ and that $fA$ is measurable so that $\operatorname{\mathbb{P}}[X\in fA]=\operatorname{\mathbb{P}}[fX\in fA]$. Then for $K>0$, arguments as above imply
$\operatorname{\mathbb{P}}[f^{-K}X\in A]=\operatorname{\mathbb{P}}[f^{-1}X\in f^{-(K-1)}A]=\operatorname{\mathbb{P}}[X\in f^{-(K-1)}A]=\operatorname{\mathbb{P}}[f^{-(K-1)}X\in A].$ (9)
Repeating this step, we obtain $\operatorname{\mathbb{P}}[f^{-K}X\in A]=\operatorname{\mathbb{P}}[X\in A]$, as required. ∎
Proof of Proposition 2.
By Proposition 1, $\Omega_{0}$ satisfies the randomization hypothesis if and only if there exists a bijective $f$ such that $\Omega_{0}\subseteq\Omega_{f}\subset\Omega_{\text{discrete}}$, where we drop measurability since all functions are measurable over discrete sets. Expanding the form of $\Omega_{f}$ using (6), we have that
$\Omega_{f}=\bigcap_{x\in\mathcal{A}^{n}}\left\{p\in\Omega_{\text{discrete}}:\prod_{j=1}^{n}p(x_{j})=\prod_{j=1}^{n}p(y_{j})\ \text{s.t.}\ y=f(x)\right\}.$ (10)
Then any $f$ such that $\Omega_{0}\subseteq\Omega_{f}\subset\Omega_{\text{discrete}}$ must have some $x^{\star}\in\mathcal{A}^{n}$ and $y^{\star}=f(x^{\star})$ such that
$\Omega_{0}\subseteq\underbrace{\left\{p\in\Omega_{\text{discrete}}:\prod_{j=1}^{n}p(x^{\star}_{j})=\prod_{j=1}^{n}p(y^{\star}_{j})\right\}}_{\equiv\Omega^{\star}}\subset\Omega_{\text{discrete}}.$ (11)
If $x^{\star}$ and $y^{\star}$ are permutations of each other so that $y^{\star}=(x^{\star}_{\sigma(1)},\dots,x^{\star}_{\sigma(n)})$ for some permutation $\sigma\in S_{n}$, then $\prod_{i=1}^{n}p(x^{\star}_{j})=\prod_{j=1}^{n}p(y^{\star}_{j})$ for all $p\in\Omega_{\text{discrete}}$, so $\Omega^{\star}=\Omega_{\text{discrete}}$. It thus must be that these two vectors are not permutations of each other. The proof is completed by showing that when this is not the case, there is some $p\in\Omega_{\text{discrete}}$ that does not belong to $\Omega^{\star}$. We can construct such a $p$ by noting that there is at least some element $\alpha\in\mathcal{A}$ that appears more times in $x^{\star}$ than in $y^{\star}$, and we can then define $p$ as placing mass $\epsilon_{0}$ on all values except $\alpha$ and mass $\epsilon_{1}\neq\epsilon_{0}$ on $\alpha$, so that $\prod_{j=1}^{n}p(x^{\star}_{j})\neq\prod_{j=1}^{n}p(y^{\star}_{j})$, as required. ∎
Proof of Proposition 3.
Define $c(\alpha;x)$ as the number of times the value $\alpha$ shows up in the vector $x$. To show that $\Omega_{0}=\{p\in\Omega_{\text{discrete}}:\operatorname{\mathbb{E}}_{p}[X_{i}^{t}]=\beta\}$ does not satisfy the randomization hypothesis, it suffices to show that for all $x,y\in\mathcal{A}^{n}$ with $c(\alpha;x)>c(\alpha;y)$ for some $\alpha\in\mathcal{A}$, there exists some $p\in\Omega_{0}$ such that $\prod_{j=1}^{n}p(x_{j})\neq\prod_{j=1}^{n}p(y_{j})$.
Let $x,y\in\mathcal{A}^{n}$ be arbitrary such that $c(\alpha;x)>c(\alpha;y)$ for some $\alpha\in\mathcal{A}$.
First, consider the case where $t$ is odd. For simplicity, let us assume that $\alpha_{1}<\alpha_{2}<\cdots<\alpha_{K}$. For $x,y\in\mathcal{A}^{n}$, we define $\beta_{i}(x,y)=c(\alpha_{i};x)-c(\alpha_{i};y)$ and notice that $\sum_{i=1}^{K}\beta_{i}(x,y)=0$. Let’s consider the following two cases:
  1. 1.
    There exists some $1\leq i\leq K-1$ such that $\alpha_{i}^{t}<\beta<\alpha^{t}_{i+1}$. Consider the following subcases:
    1. (a)
      There exists some $i_{1}\leq i$ and $i_{2}\geq i+1$ such that $|\beta_{i_{1}}(x,y)|\neq|\beta_{i_{2}}(x,y)|$. Let $p(\alpha_{j})=\epsilon$ for all $j\neq i_{1},i_{2}$, $p(\alpha_{i_{1}})=q$, and $p(\alpha_{i_{2}})=1-(K-2)\epsilon-q$. By letting $\epsilon$ to be arbitrarily small, we can ensure $\prod_{j=1}^{K}p(x_{j})\neq\prod_{j=1}^{K}p(y_{j})$ as one of the products will have more $\epsilon$ factors. We then adjust $q$ such that $\sum_{j=1}^{K}\alpha_{j}^{t}p(\alpha_{j})=\beta$.
    2. (b)
      For all $1\leq j\leq K$, $|\beta_{j}(x,y)|=C>0$ for some positive integer $C$. Because $|\mathcal{A}|>2$, we can choose $\alpha_{i_{1}}$ and $\alpha_{i_{2}}$ such that $|\alpha^{t}_{i_{1}}-\beta|\neq|\alpha^{t}_{i_{2}}-\beta|$. Then let $p(\alpha_{j})=\epsilon$ for $j\neq i_{1},i_{2}$, $p(\alpha_{i_{1}})=q$, and $p(\alpha_{i_{2}})=1-(K-2)\epsilon-q$. The argument above still applies because in order for $\sum_{j=1}^{K}\alpha_{j}^{t}p(\alpha_{j})=\beta$ to hold, we need $q\approx\frac{\beta-\alpha_{i_{2}}^{t}}{\alpha_{i_{1}}^{t}-\alpha_{i_{2}}^{t}}.$ Since $|\alpha^{t}_{i_{1}}-\beta|\neq|\alpha^{t}_{i_{2}}-\beta|$, $1-(K-2)\epsilon-q\neq q$. Although $\prod_{j=1}^{K}p(x_{j})$ and $\prod_{j=1}^{K}p(y_{j})$ may have the same number of $\epsilon$ factors, $1-(K-2)\epsilon-q\neq q$ prevents the two product to be equal.
  2. 2.
    There exists some $2\leq i\leq K-1$ such that $\beta=\alpha_{i}^{t}$. If $\beta_{i}(x,y)\neq 0$, we can just let $p(\alpha_{i})=1-\sum_{j=1,j\neq i}^{K}\epsilon_{j}$ and $p(\alpha_{j})=\epsilon_{j}$ for $j\neq i$. By letting $\epsilon_{j}$ to be arbitrarily small, we can still ensure that $\sum_{k=1}^{K}\alpha_{k}^{t}p(\alpha_{k})=\beta$ and $\prod_{k=1}^{K}p(x_{k})\neq\prod_{k=1}^{K}p(y_{k})$. For $x,y\in\mathcal{A}^{n}$ such that $\beta_{i}(x,y)=0$, we just repeat the argument for case 1 without choosing $i_{1}$ or $i_{2}$ to be $i$ and here we need $|\mathcal{A}|>3$.
When $t$ is even, we can still relabel $\alpha_{i}$ such that $\alpha_{1}^{t}\leq\alpha_{2}^{t}\leq\cdots\leq\alpha_{K}^{t}.$ With the requirement that $|\mathcal{A}|>4$, we can easily generalizes the argument for odd moments except for the problematic case where $\beta=\alpha_{i}^{t}$, $\beta_{i}(x,y)=0$, $|\beta_{j}(x,y)|=C>0$ for all $1\leq j\leq K$ and $j\neq i$, and $|\alpha_{j}^{t}-\beta|=D>0$ for all $j\neq i$. In this case, $|\{\alpha_{i}^{t}|1\leq i\leq K\}|=3$. Since $|\mathcal{A}|>4$ and $\sum_{j=1}^{K}\beta_{j}(x,y)=0$, we can only have $\alpha_{1}^{t}=\alpha_{2}^{t}$ and $\alpha_{K-1}^{t}=\alpha_{K}^{t}$. Then we can let $p(\alpha_{1})=p(\alpha_{2})=1/4-1/4\epsilon$, $p(\alpha_{K})=1/2-1/2\epsilon-\delta$ and $p(\alpha_{K-1})=\delta$ and $p(\alpha_{i})=\epsilon$. By letting $\epsilon$ and $\delta$ to be arbitrarily small, we can repeat the same argument as above.
When $|\mathcal{A}|=4$, the set $\mathcal{A}=\{-1,0,1,\sqrt{2}\}$ with the null of the second moment equaling one is equivalent to the null that 0 and $\sqrt{2}$ have equal mass. A randomization test can be constructed for this latter null by permuting these two numbers (see Example 1 of Appendix A).
Proof of Proposition 4.
Let $x,y\in\mathcal{A}^{n}$ be arbitrary such that $c(\alpha;x)>c(\alpha;y)$ for some $\alpha\in\mathcal{A}$ and let $\beta_{i}(x,y)=c(\alpha_{i};x)-c(\alpha_{i};y)$. We know that there exists some $1\leq i\leq K-1$ such that $\alpha_{i}\leq q<\alpha_{i+1}$. One of the following two cases must occur:
  1. 1.
    There exists some $i_{1}\leq i$ and $i_{2}>i$ such that $|\beta_{i_{1}}(x,y)|\neq|\beta_{i_{2}}(x,y)|$. Then we let $p(\alpha_{j})=\epsilon$ for $j\neq{i_{1}},{i_{2}}$, $p(\alpha_{i_{1}})=p-(i-1)\epsilon$, and $p(\alpha_{i_{2}})=1-(K-i-1)\epsilon-p$. One of the $\prod_{j=1}^{K}p(x_{j})$ and $\prod_{j=1}^{K}p(y_{j})$ will have more $\epsilon$ factors and hence are not the same.
  2. 2.
    For all $1\leq j\leq K$, $|\beta_{j}(x,y)|=C>0$ for some constant $C$. Because $|\mathcal{A}|>2$, without loss of generality, we can assume that there exists $i_{1},i_{2},i_{3}$ such that $i_{1}<i_{2}\leq i<i_{3}$. Then we let $p(\alpha_{j})=\epsilon$ for $j\neq{i_{1}},{i_{2}},{i_{3}}$, $p(\alpha_{i_{1}})=\frac{1}{2}p$, $p(\alpha_{i_{2}})=\frac{1}{2}p-(i-2)\epsilon$, and $p(\alpha_{i_{3}})=1-(K-i-1)\epsilon-p$. The argument above still applies because one of the $\prod_{j=1}^{K}p(x_{j})$ and $\prod_{j=1}^{K}p(y_{j})$ will have more $\epsilon$ factors.
Proof of Proposition 5.
Since $\mathbb{G}$ is a group of linear maps, each element is an invertible $n\times n$ matrix $A=\{a_{ij}\}_{ij}$. Suppose there exists $i,i^{\prime}$ and some $j^{\prime}$ such that $a_{ij^{\prime}},a_{i^{\prime}j^{\prime}}\neq 0$. If $\Omega_{A}$ is not empty, then
$X_{i}=\sum_{j}a_{ij}X_{j}$ (12)
$X_{i^{\prime}}=\sum_{j}a_{i^{\prime}j}X_{j},$ (13)
where $\{X_{j}\}_{j}$ are independent and $X_{i},X_{i^{\prime}}$ are also independent (otherwise $AX\not\stackrel{{\scriptstyle d}}{{=}}X$). We can thus apply the Darmois-Skitovich Theorem (for a reference, see Chapter 3 of Kagan et al. 1973), which implies that since $a_{ij^{\prime}}a_{i^{\prime}j^{\prime}}\neq 0$, $X_{j^{\prime}}$ is Gaussian. It follows that $X_{1},\dots,X_{n}$ are i.i.d. Gaussian, so that $\Omega_{A}$ is a subset of Gaussian distributions.
Alternatively, suppose there does not exist $i,i^{\prime}$ and some $j^{\prime}$ such that $a_{ij^{\prime}},a_{i^{\prime}j^{\prime}}\neq 0$. Then for any $j$, $\{a_{1j},\dots,a_{nj}\}$ is zero for all entries except for some index $i_{j}$. Since $A$ is invertible, for each row, at least one entry must be non-zero. Since each column has exactly one non-zero entry and since each row must have one non-zero entry, it must be that
$\begin{bmatrix}d_{1}&0&\dots&0\\ 0&d_{2}&0&\vdots\\ \vdots&0&\ddots&0\\ 0&\dots&0&d_{n}\end{bmatrix}\sigma_{n}$ (14)
where $d_{i}\neq 0$ and $\sigma_{n}$ is a permutation of the indices of $(x_{1},\dots,x_{n})$. Then, given $i$, $X_{i}=d_{j}X_{j}$ for some $j$. Then $\operatorname{Var}[X_{i}]=d_{j}^{2}\operatorname{Var}[X_{i}]$, and thus for $\Omega_{A}$ to be non-empty, it must be that $d_{j}\in\{-1,1\}$. If $d_{j}=-1$, then for $X_{i}$ with underlying distribution $P$ to lie in $\Omega_{A}$, it must be that $\operatorname{\mathbb{E}}[X_{i}]=-\operatorname{\mathbb{E}}[X_{i}]$ so that $\operatorname{\mathbb{E}}[X_{i}]=0$ and that $X_{i}=-X_{i}$, which is satisfied if and only if $P$ is symmetric around zero. Thus $\Omega_{A}$ equals the distributions that are symmetric around zero. Finally, if $d_{j}=1$, then $X_{i}=X_{i}$ is satisfied by any distribution, so that $\Omega_{A}=\Omega$. Since $\mathbb{G}$ is a collection of linear invertible matrices, $\Omega_{\mathbb{G}}$ is an intersection of these possible sets of distributions, as required. ∎
Proof of Proposition 7.
Let $X$ be a continuous random variable with probability density function $p$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]$. We will show that for any $\beta\in\operatorname{\mathbb{R}}$ we can find an $\alpha\in(0,1]$ and a probability density function $g$ on the real line with $(2k+1)$-th moment equal to $\beta$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$. Let $h$ be a probability density function supported on $\operatorname{\mathbb{R}}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=ap(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then
$\operatorname{\mathbb{E}}[X_{a}^{2k+1}]=a\operatorname{\mathbb{E}}[X^{2k+1}]+(1-a)\operatorname{\mathbb{E}}[Y^{2k+1}].$
Let $P$ be a positive number such that $P^{2k+1}>\beta$ and $Q$ be a negative number such that $Q^{2k+1}<\beta$. Let
$h(x)=\begin{cases}\mathbb{1}_{[P,P+1]}&\operatorname{\mathbb{E}}[X^{2k+1}]\leq\beta\\ \mathbb{1}_{[Q-1,Q]}&\operatorname{\mathbb{E}}[X^{2k+1}]>\beta\end{cases}$ (15)
Then,
$\operatorname{\mathbb{E}}[Y^{2k+1}]=\begin{cases}\frac{1}{2k+2}\sum_{i=0}^{2k+1}P^{i}(P+1)^{{2k+1}-i}\geq P^{2k+1}>\beta&\operatorname{\mathbb{E}}[X^{2k+1}]\leq\beta\\ \frac{1}{2k+2}\sum_{i=0}^{2k+1}(Q-1)^{i}Q^{{2k+1}-i}\leq Q^{2k+1}<\beta&\operatorname{\mathbb{E}}[X^{2k+1}]>\beta.\end{cases}$ (16)
In both cases, we can find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X_{\alpha}^{2k+1}]=\beta$.
Thus, for any $\beta\in\operatorname{\mathbb{R}}$ and $k\in\mathbb{Z}_{\geq 0}$, $\{P\in\Omega_{\text{cts}}:\operatorname{\mathbb{E}}_{P}[X_{i}^{2k+1}]=\beta\}$ is locally-dense.
We now consider the $2k$-th moment, and thus require that $\beta>0$ (else the null hypothesis is empty). Let $X$ be a continuous random variable with probability density function $p$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]$ where $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\beta^{\frac{1}{2k}}}{m+1}$. We can find an $\alpha\in(0,1]$ and a probability density function $g$ on the real line with $2k$-th moment equal to $\beta$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$. Notice that the choice of the upper bound for the interval length plays a crucial role in the definition of the locally-dense. If we remove that condition from the definition, then we can choose a probability density function $p(x)$ to be supported on $\operatorname{\mathbb{R}}\setminus[-2\beta^{\frac{1}{2k}},2\beta^{\frac{1}{2k}}]$. In this case, whatever probability density function $g$ and $\alpha\in(0,1]$ we choose such that $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$, the $2k$-th moment of the random variable with density $g$ is strictly greater than $\beta$.
Again, let $h$ be a probability density function supported on $\operatorname{\mathbb{R}}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=ap(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then
$\operatorname{\mathbb{E}}[X_{a}^{2k}]=a\operatorname{\mathbb{E}}[X^{2k}]+(1-a)\operatorname{\mathbb{E}}[Y^{2k}].$
If $\operatorname{\mathbb{E}}[X^{2k}]\leq\beta$, let $P$ be a positive number such that $P^{2k}>\beta$ and $P,P+1]\subset(\operatorname{\mathbb{R}}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}])$ and let
$h(x)=\mathbb{1}_{[P,P+1]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k}]\geq P^{2k}>\beta,$
we can easily find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k}_{\alpha}]=\beta$.
On the other hand, if $\operatorname{\mathbb{E}}[X^{2k}]>\beta$, given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\beta^{\frac{1}{2k}}}{m+1}$, we can find an interval $x,y]\subset[0,\beta^{\frac{1}{2k}}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Let
$h(x)=\frac{1}{y-x}\mathbb{1}_{[x,y]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k}]\leq y^{2k}\leq\beta.$
We can then find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k}_{\alpha}]=\beta$.
Thus, for any $\beta\in\operatorname{\mathbb{R}}_{>0}$ and $k\in\mathbb{Z}_{\geq 1}$, $\{P\in\Omega_{\text{cts}}:\operatorname{\mathbb{E}}_{P}[X_{i}^{2k}]=\beta\}$ is locally-dense. ∎
Proof of Proposition 8.
Let $X$ be a continuous random variable with probability density function $f$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]$. We will show that for any $p\in(0,1)$ and $q\in\operatorname{\mathbb{R}}$, we can find an $\alpha\in(0,1]$ and a probability density function $g$ on the real line with the $p$-th quantile equal to $q$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha f(x)$.
Let $h$ be a probability density function supported on $\operatorname{\mathbb{R}}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=af(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then,
$\operatorname{\mathbb{P}}[X_{a}\leq q]=a\operatorname{\mathbb{P}}[X\leq q]+(1-a)\operatorname{\mathbb{P}}[Y\leq q].$
There exists a large integer $N$ such that $x_{1},y_{1}],\cdots,[x_{m},y_{m}],q\subset[-N+1,N-1]$. Choose $\alpha>0$ small enough such that $\alpha\operatorname{\mathbb{P}}[X\leq q]<p$. In order for $\operatorname{\mathbb{P}}[X_{\alpha}\leq q]=p$, we need $\operatorname{\mathbb{P}}[Y\leq q]=\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}.$ This can be achieved by letting $h(x)=\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}\mathbb{1}_{[-N-1,-N]}+(1-\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha})\mathbb{1}_{[N,N+1]}$.
B.2 Proof of Proposition 6
We prove Proposition 6 by making use of Proposition 1 and showing that if $\Omega_{0}$ is locally-dense, then there is no function that is invariant to all distributions in $\Omega_{0}$ that is also not invariant to all distributions in $\Omega$. To do so, we partition the space of functions into three classes of functions. The first are ‘approximate permutation’ maps that behave like permutation maps. We will show that these functions are invariant to all continuous distributions, so cannot be used to construct a randomization test. The other two classes of functions will satisfy properties that allow us to construct, for any such function, a distribution in $\Omega_{0}$ that is not invariant to the function, which would yield the result.
B.2.1 Definitions
We begin by defining some notations and concepts. For a subset $A\subseteq\operatorname{\mathbb{R}}^{n}$, we let $m(A)$ denote the Lebesgue measure of the set. We let $\sigma_{1},\cdots,\sigma_{n!}$ be the distinct $n!$ permutations of coordinates. Let $A_{i}=\{\sigma_{i}(x):x\in A\}$ be the image of $A$ under the permutation map $\sigma_{i}$. Let $\bar{A}=\cup_{i=1}^{n!}A_{i}$ be the union of all the images of $A$ under permutation maps.
Definition 3.
We call a bijective function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$ mass-preserving if for any measurable set $A\subset\operatorname{\mathbb{R}}^{n}$, $m(A)=0$ if and only if $m(f(A))=0$.
Definition 4.
A bijective function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$ is orbit-preserving if for any measurable set $A\subset\operatorname{\mathbb{R}}^{n}$, $m(f(A)\setminus\bar{A})=0$.
Definition 5.
A bijective function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$ is box-orbit-preserving if for any box $A=\prod_{i=1}^{n}[a_{i},b_{i}]\subset\operatorname{\mathbb{R}}^{n}$, $m(f(A)\setminus\bar{A})=0$.
Definition 6.
We call a bijective function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$ an approximate permutation map if $f$ is box-orbit-preserving and mass-preserving.
Notice that both orbit-preserving and box-orbit-preserving properties imply that for any measurable set $A\subset\operatorname{\mathbb{R}}^{n}$, the image of $A$ under the approximate permutation maps are “contained” in the image of $A$ under the actual permutation maps, not the other way around. However, it suffices to require only box-orbit-preserving and mass-preserving properties to define approximate permutation maps because these properties apply to all the measurable sets $A$ and this turns out to be exactly as strong as we need.
B.2.2 Proof strategy
A function is either an approximate permutation map, or not mass-preserving, or mass-preserving but not box-orbit-preserving. The proof of Proposition 6 will follow by showing that functions that are approximate permutation maps are invariant to all distributions, while functions in the latter two classes must always fail to be invariant to some distribution in $\Omega_{0}$ when it is locally-dense. We consider each space of functions separately in the following series of lemmas, which will immediately obtain the main result. The proofs of the Lemmas are provided in Appendix B.2.3 to not disrupt the flow of the arguments here.
Lemma B.1.
A box-orbit-preserving map is orbit-preserving.
Lemma B.2.
An approximate permutation map is invariant to all i.i.d. continuous distributions.
We next consider functions that are not approximate permutation maps. These functions are either not mass-preserving or mass-preserving but not box-orbit-preserving. For such functions, we will essentially find several small intervals on which the map $f$ acts very differently from the permutation maps. On those intervals, we can define distributions in $\Omega_{0}$ that are not preserved by $f$. Outside those intervals, we will define distributions based on the locally-dense property. Consequently, when we combine these distributions, we will have a $f$-specific distribution that is not invariant to $f$ and belongs to $\Omega_{0}$.
Lemma B.3.
For any function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$ that is not mass-preserving, there exists a random vector $X=(X_{1},\dots,X_{n})$, $X_{i}$ drawn i.i.d. from distribution $P\in\Omega_{0}$ where $\Omega_{0}$ is locally-dense such that $X$ is not invariant under $f$.
Lemma B.4.
For any mass-preserving and not box-orbit-preserving function $f:\operatorname{\mathbb{R}}^{n}\rightarrow\operatorname{\mathbb{R}}^{n}$, there exists a random vector $X=(X_{1},\dots,X_{n})$, $X_{i}$ drawn i.i.d. from distribution $P\in\Omega_{0}$ where $\Omega_{0}$ is locally-dense such that $X$ is not invariant under $f$.
The proof of Proposition 6 follows immediately from the above lemmas.
Proof of Proposition 6.
By definition, a function is either an approximate permutation map, not mass-preserving, or mass-preserving but not box-orbit-preserving. If it is an approximate permutation map, Lemma B.1 shows it is invariant to all distributions. If it is not, Lemmas B.2-B.4 show it is not invariant to all distributions in $\Omega_{0}$ when it is locally-dense. The result then follows by Proposition 1. ∎
B.2.3 Proof of lemmas
Proof of Lemma B.1.
Assume for contradiction that $f$ is box-orbit-preserving and not orbit-preserving. Then there exists a measurable set $A\subset\operatorname{\mathbb{R}}^{n}$ such that $m(f(A)\setminus\bar{A})=c>0$. Let $B=f(A)\setminus\bar{A}$ and note that $B\cap\overline{f^{-1}(B)}=\emptyset$. Since $f$ is mass-preserving, $m(f^{-1}(B))>0$. We can find a finite number of pairwise disjoint boxes $D_{1},\cdots,D_{k}$ where $D_{j}=\prod_{i=1}^{n}[a^{j}_{i},b^{j}_{i}]$ such that $f^{-1}(B)\subset\cup_{j=1}^{k}D_{j}$ and $m(\cup_{j=1}^{k}D\setminus f^{-1}(B))<\frac{c}{2n!}$. Let $D=\cup_{j=1}^{k}D_{j}$.
$m(f(D)\setminus\bar{D})\geq m(B\setminus\bar{D})=m(B\setminus(\overline{D\setminus f^{-1}(B)}))\geq m(B)-n!m(D\setminus f^{-1}(B))\geq\frac{c}{2}>0.$
Thus, there exists at least one $D_{j}$ such that $m(f(D_{j})\setminus\bar{D_{j}})>0$. ∎
Proof of Lemma B.2.
First, for any measurable set $A\subset\operatorname{\mathbb{R}}^{n}$, we can find a disjoint partition of $f(A)=\bigsqcup_{i=1}^{n!}B_{i}\bigsqcup C$ such that $B_{i}\subset\sigma_{i}(A)$, $C\subset\bar{A}^{c}$. By Lemma B.1, we know that an approximate permutation map is also orbit-preserving. Thus, $m(f(A)\setminus\bar{A})=0$ and $m(C)=0$. Let $A_{i}=f^{-1}(B_{i}).$ We would like to show that $\sigma_{i}(A_{i})=B_{i}$ almost everywhere, i.e., $m(\sigma_{i}(A_{i})\triangle B_{i})=0$, where $\sigma_{i}(A_{i})\triangle B_{i}$ denotes the symmetric difference between two sets $A_{i}$ and $B_{i}$. Then, it follows
$\begin{split}\operatorname{\mathbb{P}}[X\in f(A)]=\sum_{i=1}^{n!}\operatorname{\mathbb{P}}[X\in B_{i}]=\sum_{i=1}^{n!}\operatorname{\mathbb{P}}[X\in\sigma_{i}(A_{i})]=\sum_{i=1}^{n!}\operatorname{\mathbb{P}}[X\in A_{i}]=\operatorname{\mathbb{P}}[X\in A].\end{split}$
Now to show that $m(\sigma_{i}(A_{i})\triangle B_{i})=0$, we need to show $m(\sigma_{i}(A_{i})\setminus B_{i})=0$ and $m(B_{i}\setminus\sigma_{i}(A_{i}))=0$. The latter one is trivial as $m(B_{i}\setminus\sigma_{i}(A_{i}))=m(f(A_{i})\setminus\sigma_{i}(A_{i}))=0$ follows from the condition given. It is only left to show that $m(\sigma_{i}(A_{i})\setminus B_{i})=m(\sigma_{i}(A_{i})\setminus f(A_{i}))=0$.
Assume for contradiction that there exists an $i$ where $m(\sigma_{i}(A_{i})\setminus f(A_{i}))>0$. Together with $m(f(A_{i})\setminus\sigma_{i}(A_{i}))=0$, it implies that $m(A_{i})=m(\sigma_{i}(A_{i}))>m(f(A_{i}))$. Since we can decompose $\bar{A}$ in terms of disjoint unions of $A_{j}$ where $m(A_{j})\geq m(f(A_{j}))$, it implies that $m(\bar{A})>m(f(\bar{A}))$ and $m(\bar{A}\setminus f(\bar{A}))>0$.
However, let $C=\bar{A}\setminus f(\bar{A})$. Since $\bar{C}\subset\bar{A}$ and $f^{-1}(C)\cap\bar{A}=\emptyset$, we have $f^{-1}(C)\cap\bar{C}=\emptyset$. Let $D=f^{-1}(C)$. Note that $D\cap\bar{C}=\emptyset$ implies that $\bar{D}\cap\bar{C}=\emptyset$. Then, $m(f(D)\setminus\bar{D})=m(C\setminus\bar{D})=m(C)>0$. This is a contradiction. ∎
Proof of Lemma B.3.
Because $f$ is not mass-preserving, there exists a measurable set $A\subset\operatorname{\mathbb{R}}^{n}$ such that $m(A)=0$ and $m(f(A))>0$. Thus, we can always find an arbitrarily small closed box $C=\prod_{i=1}^{n}[x_{i},y_{i}]$ where $x_{i},y_{i}]$ and $x_{j},y_{j}]$ are either disjoint or equal such that $m(C\cap f(A))>0$.
Consider the following probability density function
$p(x)=\sum_{i=1}^{n}\frac{\mathbb{1}[x_{i},y_{i}]}{n|y_{i}-x_{i}|}.$
By the definition of locally-dense, we can find an $\alpha\in(0,1]$ and a distribution $G\in\Omega_{0}$ with density function $g$ such that $g(x)|_{\bigcup_{i=1}^{n}[x_{i},y_{i}]}=\alpha p(x)$. Let $X_{1},\cdots,X_{n}$ be random variable with probability density $g$.
Let $X=(X_{1},\cdots,X_{n})$. Then
$\operatorname{\mathbb{P}}[X\in C\cap f(A)]=\frac{m(C\cap f(A))}{m(C)}\operatorname{\mathbb{P}}[X\in C]>0=\operatorname{\mathbb{P}}[X\in f^{-1}(C)\cap A],$
where the last equality is due to the fact that $m(A)=0$. Clearly, $X$ is not invariant under $f$. ∎
Proof of Lemma B.4.
Let $n$ be a positive integer. A partition of $n$ is an sequence $\lambda=(\lambda_{1},\lambda_{2},\cdots)$ of non-negative integers in decreasing order
$\lambda_{1}\geq\lambda_{2}\geq\cdots$
such that $\lambda_{1}+\lambda_{2}+\cdots=n$. Let $P_{n}$ denotes the set of all partitions of $n$. We will use $L_{n}$ to denote the reverse lexicographic ordering on the set $P_{n}$ where $L_{n}$ is the subset of $P_{n}\times P_{n}$ consisting of all $(\lambda,\mu)$ such that either $\lambda=\mu$ or the first non-vanishing difference $\lambda_{i}-\mu_{i}$ is positive. $L_{n}$ is a total ordering. We prove the lemma by induction on the reverse lexicographic ordering on partitions of $n$.
Since $f$ is not box-orbit-preserving, there exists a box $A=\prod_{i=1}^{n}[u_{i},v_{i}]$ such that
$m(f(A)\setminus\bar{A})>0.$
Partition $A$ into finitely many disjoint smaller boxes of the form $\prod_{i=1}^{n}[x_{i},y_{i}]$, where each interval $x_{i},y_{i}]$ is either disjoint from or equal to another. Since $m(f(A)\setminus\bar{A})>0$, there exists at least one such box $B$ with
$m(f(B)\setminus\bar{B})>0.$
Each such box $B=\prod_{i=1}^{n}[x_{i},y_{i}]$ induces a partition $\lambda\in P_{n}$ by recording the multiplicities of the intervals $x_{i},y_{i}]$. We proceed by induction on $P_{n}$ under the reverse lexicographic order.
Base case: Suppose $\lambda=(n)$, so $B=[x_{1},y_{1}]^{n}$. Let $C=f(B)\setminus\bar{B}$. Since $m(C)>0$ and $f$ is mass-preserving, $m(f^{-1}(C))>0$.
Consider the density
$p(x)=\frac{1}{|y_{1}-x_{1}|}\mathbf{1}_{[x_{1},y_{1}]}(x).$
By the locally-dense property, there exist $\alpha\in(0,1]$ and a distribution $G\in\Omega_{0}$ with density $g$ such that $g(x)=\alpha p(x)$ on $x_{1},y_{1}]$. Let $X=(X_{1},\dots,X_{n})$ where the $X_{i}$ are drawn i.i.d. from $G$. Then
$\mathbb{P}(X\in f^{-1}(C))=\frac{m(f^{-1}(C))}{m(B)}\alpha^{n}>0,\quad\text{and}\quad\mathbb{P}(X\in C)=0,$
so $X$ is not invariant under $f$.
Inductive step: Assume the statement holds for all partitions $\mu\succ\lambda$.
Let $B=\prod_{i=1}^{n}[x_{i},y_{i}]$ correspond to $\lambda$, and rename the intervals so that $a_{i},b_{i}]$ appears $\lambda_{i}$ times. Define the set
$S=\left\{D=\prod_{i=1}^{n}[w_{i},z_{i}]:[w_{i},z_{i}]\subset\{[a_{1},b_{1}],\dots,[a_{m},b_{m}]\}\right\}\setminus\bar{B},$
and let $S_{\leq}\subset S$ denote those boxes corresponding to partitions less than or equal to $\lambda$ and let $S_{>}\subset S$ denote those boxes corresponding to partitions larger than $\lambda$.
We consider three cases:
Case 1: Suppose there exists $D\in S_{>}$ such that $m(f(B)\cap D)>0$.
Then by the induction hypothesis, there exists a distribution $X$ (with i.i.d. coordinates from a distribution in $\Omega_{0}$) such that $X$ is not invariant under $f$. This is because $\bar{B}\cap\bar{D}=\emptyset$ and $m(f^{-1}(D)\setminus\bar{D})\geq m(f^{-1}(D)\cap B)>0$.
Case 2: Suppose there exists $D\in S_{\leq}$ such that $m(f(B)\cap D)>0$. Let $\mu_{i}$ denotes the number of times $a_{i},b_{i}]$ appears in $D$. Note that $(\mu_{1},\cdots,\mu_{m})$ might not be a partition, but the first non-vanishing difference of $\lambda_{i}-\mu_{i}$ must be positive. Let $j$ be the smallest integer such that $\lambda_{j}>\mu_{j}$.
Let $(\mu_{1},\dots,\mu_{m})$ be the corresponding multiplicities for $D$, and let $j$ be the smallest index such that $\lambda_{j}>\mu_{j}$. Define the density
$p(x)=\sum_{i=1}^{m}(\epsilon_{i-1}-\epsilon_{i})\frac{\mathbf{1}_{[a_{i},b_{i}]}(x)}{|b_{i}-a_{i}|},$
where $\epsilon_{0}=1$, $\epsilon_{m}=0$, and the $\epsilon_{i}$ are chosen later.
By the locally-dense property, there exist $\alpha\in(0,1]$ and $G\in\Omega_{0}$ with density $g$ such that $g(x)=\alpha p(x)$ on $\bigcup_{i}[a_{i},b_{i}]$. Let $X=(X_{1},\dots,X_{n})$ be i.i.d. samples from $G$.
Let $E=f(B)\cap D$. Then $m(f^{-1}(E))>0$, and
$\mathbb{P}(X\in E)\leq\operatorname{\mathbb{P}}(X\in D)=\alpha^{n}\prod_{i=1}^{m}(\epsilon_{i-1}-\epsilon_{i})^{\mu_{i}},$
$\mathbb{P}(X\in f^{-1}(E))=\frac{m(f^{-1}(E))}{m(B)}\alpha^{n}\prod_{i=1}^{m}(\epsilon_{i-1}-\epsilon_{i})^{\lambda_{i}}.$
We choose $\epsilon_{i-1}-\epsilon_{i}=\delta$ small for all $i>j$, and choose $\epsilon_{j-1}-\epsilon_{j}$ so that
$\epsilon_{j-1}-\epsilon_{j}>\frac{m(B)}{m(f^{-1}(E))}\delta.$
Thus,
$\frac{\mathbb{P}(X\in f^{-1}(E))}{\mathbb{P}(X\in E)}\geq\left(\frac{\epsilon_{j-1}-\epsilon_{j}}{\delta}\right)^{\lambda_{j}-\mu_{j}}\frac{m(f^{-1}(E))}{m(B)}>1,$
so $X$ is not invariant under $f$.
Case 3: Suppose $m\left(f(B)\cap S^{c}\cap\bar{B}^{c}\right)>0.$
Define
$p(x)=\sum_{i=1}^{m}\frac{1}{m|a_{i}-b_{i}|}\mathbf{1}_{[a_{i},b_{i}]}(x),$
and apply the locally-dense property to find $\alpha\in(0,1]$ and $G\in\Omega_{0}$ such that $g(x)=\alpha p(x)$ on $\bigcup_{i}[a_{i},b_{i}]$. Let $X=(X_{1},\dots,X_{n})$ be i.i.d. samples from $G$. Let $F=f(B)\cap S^{c}\cap\bar{B}^{c}$.
$\mathbb{P}(X\in F)\leq\operatorname{\mathbb{P}}(X\in S^{c}\cap\bar{B}^{c})=0,$
but
$\mathbb{P}(X\in f^{-1}(F))=\frac{m(f^{-1}(F))}{m(B)}\mathbb{P}(X\in B)>0,$
so again $X$ is not invariant under $f$. This concludes the proof. ∎
B.3 Statements and proofs of claims in Remark 4
Proposition B.9.
The statement in Proposition 7 holds if we replace $\Omega_{\text{cts}}(\mathbb{R})$ with $\Omega_{\text{cts}}([b_{0},b_{1}])$ for some interval $b_{0},b_{1}]$ and $\beta$ in the support. The same holds with Proposition 8 for any $q\in(b_{0},b_{1})$.
Proof of Proposition B.9.
We first consider the statement in Proposition 7 replacing $\Omega_{\text{cts}}(\mathbb{R})$ with $\Omega_{\text{cts}}([b_{0},b_{1}])$ for some interval $b_{0},b_{1}]$. We will separately consider the cases where $\beta$ is odd and even.
For the case where the moment order is odd, consider any positive integer $k$, and consider $\beta$ such that $b_{0}^{2k+1}<\beta<b_{1}^{2k+1}$. For any $m$, we will show that the null is locally-dense by considering intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset\mathcal{B}$ where
$\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\min\{b_{1}-\beta^{\frac{1}{2k+1}},\beta^{\frac{1}{2k+1}}-b_{0}\}}{m+1}.$
Consider any $X$ with probability density $p$ on such the intervals. We want to show that we can find an $\alpha\in(0,1]$ and a probability density function $g$ on $\mathcal{B}$ with $2k+1$-th moment equal to $\beta$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$. To do so, let $h$ be a probability density function supported on $\mathcal{B}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=ap(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then
$\operatorname{\mathbb{E}}[X_{a}^{2k+1}]=a\operatorname{\mathbb{E}}[X^{2k+1}]+(1-a)\operatorname{\mathbb{E}}[Y^{2k+1}].$
If $\operatorname{\mathbb{E}}[X^{2k+1}]\geq\beta$, given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\beta^{\frac{1}{2k+1}}-b_{0}}{m+1}$, we can find an interval $x,y]\subset[b_{0},\beta^{\frac{1}{2k+1}}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Let
$h(x)=\frac{1}{y-x}\mathbb{1}_{[x,y]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k+1}]\leq y^{2k+1}\leq\beta.$
We can then find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k+1}_{\alpha}]=\beta$.
Similarly, if $\operatorname{\mathbb{E}}[X^{2k+1}]\leq\beta$, given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{b_{1}-\beta^{\frac{1}{2k+1}}}{m+1}$, we can find an interval $x,y]\subset[\beta^{\frac{1}{2k+1}},b_{1}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Let
$h(x)=\frac{1}{y-x}\mathbb{1}_{[x,y]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k+1}]\geq x^{2k+1}\geq\beta.$
We can then find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k+1}_{\alpha}]=\beta$.
For the case where the moment order is even, we will show that for any real number $0<\beta<\max\{b_{0}^{2k},b_{1}^{2k}\}$ and any continuous random variable $X$ with probability density function $p$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset\mathcal{B}$ where
$\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\min\{\max\{|b_{0}|,|b_{1}|\}-\beta^{\frac{1}{2k}},\beta^{\frac{1}{2k}}\}}{m+1},$
we can find an $\alpha\in(0,1]$ and a probability density function $g$ on $\mathcal{B}$ with $2k$-th moment equal to $\beta$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$.
To show this, let us assume without loss of generality that $|b_{1}|\geq|b_{0}|$. Let $h$ be a probability density function supported on $\mathcal{B}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=ap(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then
$\operatorname{\mathbb{E}}[X_{a}^{2k}]=a\operatorname{\mathbb{E}}[X^{2k}]+(1-a)\operatorname{\mathbb{E}}[Y^{2k}].$
If $\operatorname{\mathbb{E}}[X^{2k}]\geq\beta$, given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{\beta^{\frac{1}{2k}}}{m+1}$, we can find an interval $x,y]\subset[-\beta^{\frac{1}{2k}},\beta^{\frac{1}{2k}}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Let
$h(x)=\frac{1}{y-x}\mathbb{1}_{[x,y]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k}]\leq y^{2k}<\beta.$
We can then find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k}_{\alpha}]=\beta$.
Similarly, if $\operatorname{\mathbb{E}}[X^{2k}]\leq\beta$, given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\frac{|b_{1}|-\beta^{\frac{1}{2k}}}{m+1}$, we can find an interval $x,y]\subset[\beta^{\frac{1}{2k}},\mathbb{1}\left[b_{1}>0\right]\times b_{1}-\mathbb{1}\left[b_{1}<0\right]\times b_{1}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Let
$h(x)=\frac{1}{y-x}\mathbb{1}_{[x,y]}.$
Then
$\operatorname{\mathbb{E}}[Y^{2k}]\geq x^{2k}\geq\beta.$
We can then find an $\alpha\in(0,1]$ such that $\operatorname{\mathbb{E}}[X^{2k}_{\alpha}]=\beta$.
We next consider the statement in Proposition 8 replacing $\Omega_{\text{cts}}(\mathbb{R})$ with $\Omega_{\text{cts}}([b_{0},b_{1}])$ for some interval $b_{0},b_{1}]$ and considering any $q\in(b_{0},b_{1})$. Let $X$ be a continuous random variable with probability density function $f$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]\subset[b_{0},b_{1}]$ where $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\min\{\frac{q-b_{0}}{m+1},\frac{b_{1}-q}{m+1}\}$. We will show that for any $p\in(0,1)$, there exist $\alpha\in(0,1]$ and a probability density function $g$ on $b_{0},b_{1}]$ such that $q$ is the $p$-th quantile of $g$ and $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha f(x)$.
Let $h$ be a probability density function supported on $b_{0},b_{1}]\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=af(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$. Then,
$\operatorname{\mathbb{P}}[X_{a}\leq q]=a\operatorname{\mathbb{P}}[X\leq q]+(1-a)\operatorname{\mathbb{P}}[Y\leq q].$
Given the condition $\max_{1\leq i\leq m}|y_{i}-x_{i}|<\min\{\frac{q-b_{0}}{m+1},\frac{b_{1}-q}{m+1}\}$, we can find intervals $x,y]\subset[b_{0},q]$ and $w,z]\subset(q,b_{1}]$ such that $x,y]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=[w,z]\cap\bigcup_{i=1}^{m}[x_{i},y_{i}]=\emptyset$.
Now we can choose $\alpha>0$ small enough such that $\alpha\operatorname{\mathbb{P}}[X\leq q]<p$. In order for $\operatorname{\mathbb{P}}[X_{\alpha}\leq q]=p$, we need $\operatorname{\mathbb{P}}[Y\leq q]=\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{1-\alpha}.$ This can be achieved by letting $h(x)=\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{(1-\alpha)(y-x)}\mathbb{1}_{[x,y]}+(1-\frac{p-\alpha\operatorname{\mathbb{P}}[X\leq q]}{(1-\alpha)})\frac{1}{(z-w)}\mathbb{1}_{[w,z]}$. ∎
Proposition B.10.
For any $\beta\in\mathbb{R}_{>0}$, $\Omega_{0}=\{P\in\Omega_{\text{cts}}(\mathbb{R}):\operatorname{Var}_{P}[X_{i}]=\beta\}$ is locally-dense and thus does not satisfy the randomization hypothesis.
Proof of Proposition B.10.
Let $X$ be a continuous random variable with probability density function $p$ supported on $m$ intervals $x_{1},y_{1}],\cdots,[x_{m},y_{m}]$. For any $\beta\in\operatorname{\mathbb{R}}_{>0}$, we will show that we can find an $\alpha\in(0,1]$ and a probability density function $g$ on the real line with variance $\beta$ such that $g(x)|_{\bigcup_{i=1}^{m}[x_{i},y_{i}]}=\alpha p(x)$.
Let $h$ be a probability density function supported on $\operatorname{\mathbb{R}}\setminus\bigcup_{i=1}^{m}[x_{i},y_{i}]$ and let $Y$ be the random variable with probability density $h$. For any $a\in(0,1]$, let $g_{a}(x)=ap(x)+(1-a)h(x)$. Let $X_{a}$ be the random variable with probability density $g_{a}$.
Then
$\begin{split}\operatorname{Var}[X_{a}]&=a(\operatorname{Var}[X]+\operatorname{\mathbb{E}}[X]^{2})+(1-a)(\operatorname{Var}[Y]+\operatorname{\mathbb{E}}[Y]^{2})-(a\operatorname{\mathbb{E}}[X]+(1-a)\operatorname{\mathbb{E}}[Y])^{2}\\ &=a[(1-a)\operatorname{\mathbb{E}}[X]^{2}+\operatorname{Var}[X]+(1-a)\operatorname{\mathbb{E}}[Y]^{2}-2(1-a)\operatorname{\mathbb{E}}[X]\operatorname{\mathbb{E}}[Y]]+(1-a)\operatorname{Var}[Y].\end{split}$
Choose $M>0$ such that $\bigcup_{i=1}^{m}[x_{i},y_{i}]\subset[-M,M]$. For any $\beta>0$ and $a=\frac{1}{2}$, we can find $z>M+1$ such that
$a[(1-a)\operatorname{\mathbb{E}}[X]^{2}+\operatorname{Var}[X]+(1-a)z^{2}-2(1-a)z\operatorname{\mathbb{E}}[X]]>\beta.$
We then choose $w\in(0,1)$ such that $\frac{1}{12}w^{2}<\beta$. Let $h(x)=\frac{1}{w}\mathbb{1}_{[z-\frac{1}{2}w,z+\frac{1}{2}w]}$. In this case, since $\operatorname{Var}[Y]\geq 0$, we have $\operatorname{Var}[X_{1/2}]>\beta$. On the other hand, $\operatorname{Var}[X_{0}]=\operatorname{Var}[Y]=\frac{1}{12}w^{2}<\beta$. Because $\operatorname{Var}[X_{a}]$ is a continuous in $a$, we can find an $\alpha\in(0,\frac{1}{2})$ such that $\operatorname{Var}[X_{\alpha}]=\beta$. Thus, for any $\beta\in\operatorname{\mathbb{R}}_{>0}$, $\{P\in\Omega_{\text{cts}}:\operatorname{Var}_{P}[X_{i}]=\beta\}$ is locally-dense. ∎